### v1

# Unveiling the Siren’s Song: Towards Reliable Fact-Conflicting Hallucination Detection

Abstract.
Large Language Models (LLMs), such as ChatGPT/GPT-4, have garnered widespread attention owing to their myriad of practical applications, yet their adoption has been constrained by issues of fact-conflicting hallucinations across web platforms. The assessment of factuality in text, produced by LLMs, remains inadequately explored, extending not only to the judgment of vanilla facts but also encompassing the evaluation of factual errors emerging in complex inferential tasks like multi-hop, and etc. In response, we introduce FactCHD, a fact-conflicting hallucination detection benchmark meticulously designed for LLMs. Functioning as a pivotal tool in evaluating factuality within “Query-Respons” contexts, our benchmark assimilates a large-scale dataset, encapsulating a broad spectrum of factuality patterns, such as vanilla, multi-hops, comparison, and set-operation patterns. A distinctive feature of our benchmark is its incorporation of fact-based chains of evidence, thereby facilitating comprehensive and conducive factual reasoning throughout the assessment process. We evaluate multiple LLMs, demonstrating the effectiveness of the benchmark and current methods fall short of faithfully detecting factual errors. Furthermore, we present Truth-Triangulator that synthesizes reflective considerations by tool-enhanced ChatGPT and LoRA-tuning based on Llama2, aiming to yield more credible detection through the amalgamation of predictive results and evidence [^1] The benchmark dataset and source code will be made available in https://github.com/zjunlp/FactCHD..
Fact-Conflicting, Hallucination Detection, LLMs, Benchmark
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
1. Introduction
Figure 1. Illustrating Fact-Conflicting Hallucination Detection.
There is a significant risk that LLMs (Zhao et al., 2023) may generate web content that appears credible but is factually incorrect or imprecise (Yin et al., 2023; Ji et al., 2023; Huang et al., 2023), thereby spreading misinformation on the web. This phenomenon is known as fact-conflicting hallucination (Zhang et al., 2023) and may result from flawed or outdated knowledge incorporated into the model’s parameters, as well as the language model’s shortcomings in tasks such as fact reconciliation, logical reasoning, and other sophisticated cognitive processes. This limitation not only restricts the application of generative AI in high-stakes areas such as finance, healthcare and law but also exacerbates the dissemination of erroneous information on the web community, posing a significant obstacle to the healthy development of the web.
Given their pivotal role in ensuring the security of online content, effectively detecting and mitigating fact-conflicting hallucinations stands paramount. Traditional task definitions for fact verification (Wadden et al., 2020a, b) exhibit limitations, particularly in accommodating writing tasks that users frequently navigate when interfacing with LLMs. Moreover, existing hallucination evaluation benchmarks (Li et al., 2023a; Muhlgay et al., 2023), predominantly centering on rudimentary facts and textual content, lack in-depth exploration of complex operations among facts, thus rendering their coverage against fact-conflicting hallucinations suboptimal. In light of this, we introduce an inherently rigorous, yet authentic, task scenario: fact-conflicting hallucination detection, devoid of explicit claims or evidence. For instance, in Figure 1, when confronted with a query and its generated response, evaluators, within the prescribed scenario, are impelled to harness both their intrinsic knowledge and external resources, while employing reasoning skills to render a factual judgment accompanied by an elucidative reasoning process.
In paving the way for future strides in hallucination evaluation, we dedicate our efforts towards establishing a rigorous benchmark, Fact-Conflicting Hallucination Detection (FactCHD), tailored for LLMs and encompassing a variegated array of factuality patterns, including Vanilla, Multi-hops, Comparison, and Set-Operation patterns. Drawing inspiration from the adage “to know it and to know the reason why of it” by Zhuzi, FactCHD endeavors to intertwine fact-based chains of evidence, providing cogent justifications when predicting the factual or non-factual nature of a statement. Achieving the intended goal of gathering a substantial amount of data through full-process human annotation is a time-consuming and resource-intensive task with limited scalability. Considering the high reliability of existing knowledge graphs (KGs) and textual knowledge, we advocate for a data construction strategy that leverages knowledge as facts, coupled with semi-supervised annotation, to facilitate the crafting of the aforementioned benchmark.
As illustrated in Figure 3, we engineer our benchmark through a systematic construction strategy, bifurcated into two primary sections, encompassing six meticulous steps. Initially, a subset of queries is input into several LLMs, with corresponding responses being manually annotated, thereby curating a minimal but realistic “Query-Response” dataset enriched with factuality information. Subsequently, we intertwine knowledge sub-graphs and textual knowledge, utilizing these as factual underpinnings to guide ChatGPT in the autonomous generation of contextual “Query-Response” pairs. Then human similarity assessment is employed iteratively to finetune prompts, ensuring stylistic coherence between generated and realistic factuality data. The fourth step involves automated similarity screening of the synthesized factuality data, identifying and excising highly similar samples to uphold dataset diversity. Fifthly, ChatGPT is guided by precise instructions to generate evidence chains, anchoring its judgments in the “Query-Response” contexts on solid knowledge. Concludingly, an adept team of 21 individuals meticulously filters the data, adhering to stipulated rules, thereby assuring the procurement of a high-quality dataset.
We conduct baseline experiments to assess the efficacy of several LLMs (such as Alpaca, ChatGPT, etc), utilizing our FactCHD under various settings—zero-shot, in-context learning, tuning for detect-specific expertise, and knowledge enhancement through retrieval/tools. Despite tuning and knowledge augmentation positively influencing the evaluation of fact-conflicting hallucinations, both open-source LLMs and ChatGPT encounter challenges in the precise and robust detection of factual inaccuracies. Thus, we introduce Truth-Triangulator, our innovative framework, inspired by the “Triangulation” theory (Valenza, 2016). Within this structure, tool-enhanced ChatGPT is designated as the Truth Seeker, and the detect-specific expert (based on Llama2-7B-LoRA) serves as the Truth Guardian. An additional agent, the Fact Verdict Manager, is tasked with amassing evidence from various perspectives to fortify the reliability and accuracy of the derived conclusions. Truth-Triangulator accentuates employing cross-referencing generators and search capabilities to astutely evaluate and adjudicate responses with potential factual discrepancies. Key insights are summarized as follows:
  • We devise FactCHD, a comprehensive large-scale factuality evaluation benchmark that encompasses multiple domains and a myriad of factual dimensions, complemented by interpretable fact-based chains of evidence. This benchmark is poised to be an invaluable asset, propelling future innovations in detecting fact-conflicting hallucinations.
  • We introduce a semi-supervised data construction strategy, utilizing knowledge as facts to expedite the assembly of the fact-conflicting hallucination dataset. Owing to the authenticity and extensive domain coverage of KGs, this strategy not only assures enhanced scalability but also stands out as particularly adept for large-scale applications.
  • We present Truth-Triangulator, a framework employing triangulation for truth discernment, utilizing cross-referencing generators and search tools to arbitrate responses of dubious factuality. Initial experiments substantiate the varied performance of different LLMs in identifying fact-conflicting hallucinations and affirm the superiority of our proposed approach.
2. Related Work
Hallucination in LLMs.
Despite the impressive capabilities of LLMs (Zhao et al., 2023; Chen et al., 2022b, a; Zhang et al., 2022), such as ChatGPT (OpenAI, 2022), in comprehending and carrying out user instructions, they sometimes generate misleading information with a high level of confidence. This phenomenon, known as hallucination (Ji et al., 2023; Ye et al., 2023), has been classified by (Zhang et al., 2023) into three categories: input-conflicting, context-conflicting, and fact-conflicting hallucinations. While humans can relatively easily identify the first two types, fact-conflicting hallucinations present a significant challenge as they involve the dissemination of misleading information on the web due to conflicting factual knowledge. While hallucination in vanilla natural language generation (NLG) settings has been extensively studied (Shuster et al., 2021; Creswell and Shanahan, 2022; Shuster et al., 2021; Das et al., 2023; Mallen et al., 2022), these methods primarily focus on analyzing hallucination across various NLP tasK.  Li et al. (2023a) recently introduced HaluEval, the hallucination evaluation benchmark evaluating the performance of LLMs in recognizing hallucinations. Different from HaLuEval which only analyzes ChatGPT’s ability to evaluate hallucination, we specifically focus on the evaluation of fact-conflicting hallucination by constructing an interpretable benchmark that can serve as a public platform for checking the factual errors of context with explanation, as well as designing the Truth-Triangulator framework for reliable hallucination detection.
Split #Sample Van. Multi. Comp. Set-Op.
Train 51,383 31,986 8,209 5,691 5,497
Test 6,960 4,451 1,013 706 790
Table 1. Distribution of our proposed benchmark, FactCHD
Table 2. Comparison with existing fact-checking datasets. Our FactCHD include both general and vertical domains, such as health, COVID-19, climate, science, and medicine (genes, virus and disease). Van., Mul., Com. and Set. are abbreviations for the four distinct factuality patterns in FactCHD. We utilize the following abbreviations: HG - human-curated datasets; MIX - datasets containing both human-constructed and machine-generated data; Q&R - “Query-Response” context; Consist. - classification of consistency between claims and given evidence; Fact. - classification of factuality, with labels factual and non-factual; Chain. - chains of evidence.
Datasets # Source Domains Construction Format Factuality Pattern Evaluation Expandability
Fever (Thorne et al., 2018) Text General HG Claim Van. Consist.
Climate-Fever (Diggelmann et al., 2020) Text Climate Change HG Claim Van. Consist.
Health-Fever (Sarrouti et al., 2021) Text Health HG Claim Van. Consist.
Sci-Fact (Wadden et al., 2020a) Text Scientific HG Claim Van. Consist.
CoVERT (Mohr et al., 2022) Text COVID-19 HG Claim Van. Consist.
TabFact (Chen et al., 2020) Table General HG Claim Van. Consist.
HoVer (Jiang et al., 2020) Text General HG Claim Mul. Consist.
FEVEROUS (Aly et al., 2021) Text+Table General HG Claim Van. Consist.
HaluEval (Li et al., 2023a) Text General MIX Q&R Van. Hallucination
FactCHD (ours) KGs (
) & Text
General&Vertical MIX Q&R Van.&Mul.&Com.&Set. Fact.+Chain.
Factuality Detection.
Our work is also related to previous fact verification or actuality evaluation in various NLP tasks (Thorne et al., 2018; Wadden et al., 2020b; Gupta et al., 2022; Dziri et al., 2022; Rashkin et al., 2021; Kryściński et al., 2020; Pagnoni et al., 2021; Tang et al., 2022). Going beyond a simplistic binary notion of factuality, the FRANK framework (Pagnoni et al., 2021) advocates for a nuanced comprehension of factuality and introduces a typology of factual errors. Dhingra et al. (2019) measures $n$-gram lexical entailment for table-to-text generation. Recently,  Muhlgay et al. (2023) proposes FACTOR score describes LM’s tendency to assign greater likelihood to the original fact than to any of the non-factual alternatives. FELM (Chen et al., 2023) extends actuality evaluation into math and reasoning tasks. However, these benchmarks and metrics may not align well with the naturally occurring “Query-Response” format in the context of LLMs, and they emphasize accuracy evaluation while neglecting the provision of explanatory justifications for decision-making. In contrast, our work (1) provides challenge evaluation data in “Query-Response” format for LLMs; (2) considers an additional dimension of evaluation metrics that evaluate the interpretability of detecting fact-conflicting hallucination; and (3) investigates the effectiveness of Truth-Triangulator with verifying and confirming information by cross-referencing multiple independent sources or perspectives. Table 2 shows the comparison with other relevant datasets. Moreover, factuality-based hallucination detection establishes the foundation for model editing (Yao et al., 2023; Mao et al., 2023; Li et al., 2023b; Wang et al., 2023) of LLMs.
3. FactCHD Benchmark Construction
We first define the task of fact-conflicting hallucination detection and subsequently detail the benchmark construction process in the following subsections.
Task Formulation. The detection of fact-conflicting hallucinations in LLM responses in real-world scenarios entails identifying factual inaccuracies and considering potential human-initiated requests. A detector should not only predict the factuality label accurately but also generate an explanatory justification for its binary prediction. Thus, the task of fact-conflicting hallucination detection is formally defined as: Input: A question $Q$ and response $R$ from LLMs, encompassing four types of fact-conflicting: Vanilla, Multi-hop, Comparison, and Set-Operation. Output: A sequence $A=[l,e]$, where $l$ and $e$ represent the factuality label (factual or non-factual) and explanation, respectively.
Based on the aforementioned definitions, we construct FactCHD, which comprises a comprehensive array of training samples, enriched with an additional 6,960 meticulously curated samples for evaluating fact-conflicting hallucinations generated by LLMs. Our dataset ensures an equitable balance between factual or non-factual categories, providing a sturdy platform for evaluation. Notably, FactCHD presents three distinct characteristics: (1) It encompasses patterns of diverse reasoning, including multi-hop, comparison, and set operations as shown in Table 1; (2) FactCHD adheres to realistic scenarios by supplying “Query-Response” pairs and their associated evidence to validate the information provided; (3) The benchmark is meticulously crafted, leveraging KGs for initial data construction, followed by meticulous human verification to assure quality. Furthermore, the dataset inherently allows for expansion through updates based on KGs, offering a unique advantage in maintaining its contemporaneity and expandability. Further specific instructions and case data are illustrated in Appendix A. We proceed to introduce the design principles of the fact-conflicting hallucination detection benchmark in the following parts.
Figure 2. Distribution of FactCHD across various domains, with “General-p” and “General-d” denoting prevalent domains sourced from Wikipedia and Wikidata, respectively.
Addressing hallucinations in the outputs of LLMs necessitates a meticulous examination of both hallucinatory and non-hallucinatory instances. The susceptibility of LLMs to fact-conflicting hallucinations stems from their circumscribed knowledge and suboptimal reasoning capacities (Ji et al., 2023). Accordingly, we categorize four distinct patterns of factual errors, vividly illustrated in the lower section of Figure 3: (1) vanilla pattern that deals with factual statements that can be objectively verified using established sources and are typically more straightforwardly identified; (2) multi-hops pattern denotes the process of drawing conclusions by connecting multiple pieces of facts, (3) Comparison pattern refers to the act of evaluating and comparing the relative worth, and relationships between different pieces of facts, and (4) Set-operation pattern involves manipulating and combining sets of elements using operations to analyze relationships between different facts. Subsequently, we query open-source LLMs using specific instructions that reflect the identified factuality patterns. Through manual annotation of the hallucination responses, we curate a genuine compilation of hallucinations. Our objective is to gather a substantial amount of bulk hallucination data to simulate these realistic hallucinations as demonstrative examples.
3.1. Generate “Query-Response” Contexts
Capitalize knowledge as Fact
KGs, with their extensive repositories of entities and relational data, not only champion compositional reasoning but also offer a foundational structure for factual information. Concurrently, textual knowledge, while invaluable for vanilla inference, takes on a pivotal role in offering additional and nuanced information. With this in mind, our aim steers toward exploring the capabilities of utilizing existent knowledge as a factual backbone in the semi-automated assembly of fact-conflicting hallucinations, detailed as follows:
(1) We extract KG data from Wikidata (Vrandecic and Krötzsch, 2014) and PrimeKG (Chandak et al., 2023), utilizing it as an underpinning knowledge base for generating “Query-Respons” data. By manually selecting $438$ prevalent relations and crafting subgraphs through $K$-hop walks, performed $N$ times from divergent, randomly selected starting entities, we procure a varied set of subgraphs to be employed for multi-hop reasoning, fact comparison, and set operation. Heuristic rules are applied to ensure minimal intersection and coherence among the extracted subgraphs.
(2) We adopt the text knowledge in these various fact verification datasets including FEVER (Thorne et al., 2018), Climate-Fever Diggelmann et al. (2020), Health-Fever (Sarrouti et al., 2021), COVID-FACT (Saakyan et al., 2021) and SCIFACT (Wadden et al., 2020a) to construct the data with vanilla pattern in our benchmark FactCHD. We only select the factual (supported) and non-factual (refuted) samples with their corresponding evidence and initially employ ChatGPT (GPT-3.5-turbo) to directly evaluate the claims in these datasets and select out samples where the model struggles to provide false answers.
Figure 3. Overview of the construction and pattern illustration of FactCHD.
3.2. Collect Realistic Data with Diverse Patterns
Prompt for Q&R Generation
We devise effective prompts to steer ChatGPT in generating the “Query-Response” contexts, comprising three vital elements: role description, factuality pattern, and factuality demonstration. The role description defines the system’s role and specifies the input and objective of the generation process. To ensure precise control over the type and quality of factuality-related samples, we give the explanation of corresponding factuality patterns and demonstrations to guide the model in generating “Query-Response” contexts.
3.3. Refine Prompts with Consistency
We utilize ChatGPT to generate “Query-Response” contexts, deploying manual prompts to enable its function as a role-playing agent, informed by specific demonstrative examples. Recognizing ChatGPT’s significant prompt sensitivity, we engage in an iterative refinement of them to ensure the generated data aligns closely with desired, realistic patterns, following the protocol delineated in § 3.1. Beginning with an initial fixed set of 100 examples, we select five samples for human similarity assessment, contrasting them with demonstrative examples. Employing the majority rule, we evaluate whether each context adheres to established similarity criteria. This iterative process allows for continuous prompt adjustment, assuring alignment between generated data patterns and desired aspirations.
3.4. Automated Screening for Enhanced Diversity
Aiming to improve the diversity of generated “Query-Response” contexts, we employ Sentence-BERT (SBERT)(Reimers and Gurevych, 2019) in an automated screening process, calculating average semantic similarity within contexts. This facilitates the identification and exclusion of highly similar samples, safeguarding dataset diversity. After the screening process, a total of 1,538 samples from the training set and 632 samples from the test set were removed, ensuring the final aggregate of samples is diverse. This meticulous elimination of semantically analogous entries ensures our collections exhibit varied and non-redundant queries and responses, enhancing the benchmark’s utility for evaluating diverse instances of fact-conflicting hallucination.
3.5. Generate Chain of Evidence
Our benchmark extends beyond merely identifying factual errors in “Query-Response” contexts produced by LLMs; it also necessitates the generation of coherent evidence chains to substantiate such judgments, all of which are rooted in factual knowledge and articulated by ChatGPT. Particularly, we exploit two types of knowledge sources (outlined in §3.1) as our factual foundation. ChatGPT employs either subgraph facts or textual facts to furnish comprehensive justifications when assigning appropriate labels to the “Query-Response” contexts. It’s pivotal to note that the efficacy of these explanations profoundly impacts the credibility and reliability of judgments, thereby fostering trust in predictive models. These explanatory outputs can also amplify users’ overall understanding.
3.6. Quality Filtering via Human Annotator
We devised filter rules encompassing three facets: pattern consistency, response factuality, and evidence chain logic, employing these to guide annotators through quality filtering. Our team of 21 annotators, each boasting proficient English reading skills and a minimum of undergraduate education, underwent uniform training per standardized annotation guidelines. The annotators leverage their own awareness and, when necessary, utilize search engines to ensure informed filtering decisions, discarding samples that flout established rules. Given the inherent subjectivity in defining matching quality via these rules, we divided both the annotators and samples into seven groups, each featuring three annotators reviewing the same data batch through a voting mechanism. Simultaneously judged mismatches by annotators led to sample discarding, resulting in final removal counts of 565 and 258 samples from the training and test sets, respectively.
4. Experiments
[…]
Evaluator Vanilla Multi-hops Comparison Set-Operation Average
cls. exp. cls. exp. cls. exp. cls. exp. cls. exp.
Zero-Shot
GPT-3.5-turbo 55.12 22.79 59.54 29.84 16.66 18.89 55.46 28.23 52.82 24.03
text-davinci-003 52.06 17.72 59.92 25.30 25.50 16.09 48.58 25.71 50.98 19.57
Alpaca-7B 29.66 11.72 5.20 25.60 8.88 17.95 13.08 21.37 23.10 13.66
Vicuna-7B 35.26 24.62 17.54 34.39 9.34 24.88 14.96 31.41 28.84 26.84
Llama2-7B-chat 3.57 26.78 5.49 33.87 10.53 35.25 12.61 33.27 5.77 29.41
ICL (4-shot)
[…]
Det.(tune)
[…]
Knowledge
[…]
Truth-Triangulator 80.97 47.08 75.01 64.21 66.27 55.70 80.87 65.25 78.15 52.52
4.1. Metrics and Settings
FactCls Metric
We employ the FactCls metric, denoted by the Micro F1 score, to assess binary factuality classification performance, with a distinct focus on the distribution $p(l|Q\&R)$, classifying instances as either factual or non-factual. In light of our commitment to pinpointing non-factual examples, we designate non-factual as the positive class and factual as the negative class.
ExpMatch Metric
In FactCHD, the chain of evidence integrates two elements: introductory, expository statements and a core of factual explanations. The former provides phrasing like ’Therefore, there is an incorrect conclusion in this query and response’ as a segue, while the latter delivers detailed insights, bolstering the reasoning behind factuality judgments. Given the paramount importance of aligning factual explanations over expository parts, we employ a segmented matching and weighted averaging metric, ExpMatch, amalgamating these sections for evaluative precision. Specifically, we determine $\text{Score}_{bd}$ using span-based Micro F1 to measure word overlap at the unigram level between the model-generated and reference bodies, reflecting factual core matching. Concurrently, $\text{Score}_{ht}$ is defined via the ROUGE-L score for the head-to-tail part, assessing similarity based on the longest common word subsequence between system-generated context and gold reference. Ultimately, we introduce the ExpMatch metric, a composite of $\text{Score}_{bd}$ and $\text{Score}_{ht}$, serving as the pivotal metric for appraising the coherence and quality of utilized factual chains in explanations:
(1) $\textsc{ExpMatch}=\alpha\times\text{Score}_{bd}+(1-\alpha)\times\text{Score}_{ht}.$
In our experiment, we initially assign $\alpha$ to 0.7. If a generated result lacks factual or non-factual information, indicating a prediction failure, we attribute an ExpMatch value of 0 to that particular example during metric calculation.
4.2. Empirical Examination of Baseline Strategies
This study seeks to probe the efficiency and viability of several prototypical baseline strategies in the detection of fact-conflicting hallucinations. These strategies encompass: (1) Zero-shot Learning, which assesses model performance without prior training in hallucination detection; (2) In-context Learning, employing 4-shot samples as demonstrations to prompt the model and evaluate its competence in managing context-specific hallucinations; (3) Detect-Specific Expert Model, finessing the LoRA (Hu et al., 2022) parameters through fine-tuning, utilizing our specialized training set to harness domain expertise; (4) Knowledge Enhancement, amplifying the model’s capabilities by integrating external knowledge via retrieval techniques or tool-based enhancements. Through these experiments, we aim to traverse both the potentials and limitations of these baseline methods within the scope of hallucination detection.
Evaluation Models.
We evaluate the performance of various prominent LLMs within the context of our FactCHD benchmark, primarily focusing on models accessible through the OpenAI API, such as text-davinci-003, which corresponds to the InstructGPT model trained using supervised instruction tuning (Ouyang et al., 2022), as well as GPT-3.5-turbo (ChatGPT), which represents one of the most widely recognized models. Alternatively, we explore the adoption of open-source models such as Llama2-chat (Touvron et al., 2023b), Alpaca (Taori et al., 2023) and Vicuna (Chiang et al., 2023), which are fine-tuned variants of the LLaMA (Touvron et al., 2023a).
Implementation Details.
We conduct the generation of factuality evaluation instances using the Azure OpenAI ChatGPT API [^2] https://openai.com/blog/chatgpt/. During the generation process, we employ a temperature of $1.0$ to control the diversity of generated samples, while limiting the maximum number of tokens to $2048$ to ensure concise responses. Additionally, we set the frequency penalty to zero and utilize a Top-$p$ value of $1.0$ to ensure unrestricted token selection during generation. For evaluation purposes, we set the temperature to 0.2 across all models to reduce output randomness. Specific instructions employed in experiments as detailed in Appendix B.
Figure 4. Overview Truth-Triangulator. Here we designate the Refer to caption as the “Truth Guardian” based on Llama2-7B-chat-LoRA while Refer to caption as the “Truth Seeker” based on GPT-3.5-turbo (tool) in our experiments. We want the Refer to caption “Fact Verdict Manager” to collect evidence from different viewpoints to enhance the reliability and accuracy of the obtained conclusion.
4.2.1. Zero-shot Learning Performance
The empirical results of the evaluation, presented in Table 3, reveal that LLMs with zero-shot learning still struggle with identifying implicit factuality in “Query-Response” contexts. The performance of the open-source 7B LLMs in zero-shot learning is notably poor, highlighting its deficiency in both instruction comprehension and internal knowledge representation. Even the ChatGPT exhibits limited proficiency in distinguishing between factual and non-factual “Query-Response” samples, achieving a mere 52.82% FactCls score in zero-shot, just slightly above chance level. Further investigation into the underwhelming performance reveals a notable drawback: LLMs (especially Llama2-chat) are susceptible to false negatives, resulting in reduced sensitivity when it comes to error detection.
4.2.2. In-context Learning Performance
[…]
4.2.3. Detect-Specific Expert Performance
This study explores the efficacy of tuning LLMs using LoRA to serve as domain-specific experts in identifying fact-conflicting hallucinations. Through training LLMs with our dedicated instructions and datasets, which are intrinsically focused on factuality, we aim to hone their proficiency in this particular domain. Our results highlight a notable enhancement across all open-source 7B models, wherein Llama2-chat emerges as the top-performing model amongst its peers. Post-fine-tuning with fact-conflicting hallucination data, Llama2-chat-7B models attain peak performance on our benchmark, as evidenced by a 74.73% FactCls score and a 53.71% ExpMatch score, respectively. The demonstrated success in training a scaled-down 7B model using our approach underlines the potential of employing 7B LLMs as effective hallucination detectors and paves the way for additional research and trials in this field.
4.2.4. Knowledge Enhancemant
Retrieval Enhancement
Leveraging retrieval of pertinent knowledge stands as a prevalent strategy to mitigate factual inaccuracies (Lewis et al., 2020). This study empowers LLMs with knowledge facts derived from Wikipedia, deploying TF-IDF for document retrieval and subsequently applying BM25 to pinpoint the top two most pertinent paragraphs from the fetched document. Given the 7B model’s constrained capacity to comprehend lengthy instructional texts—as unveiled in our ICL experiments—we elevate its grasp of expansive knowledge by supplying knowledge for fine-tuning the LoRA parameters. However, compared to the detect-specific expert without knowledge augmentation, the improvement achieved by combining Wikipedia retrieval with fine-tuning the LoRA parameters is relatively modest. We attribute these observations to two primary factors: our dataset spans multiple domains while Wikipedia knowledge encompasses merely a subset, and our retrieval method, being elementary and generic, occasionally yields lower-quality evidence.
Table 4. Out-of-distribution case analysis of our Truth-Triangulator.
User Contexts and Returned Evidence Detection Opinion from Both Sides Prediction with Triangulation
Query: Who was the mother of Afonso II, the third king of Portugal?
Response:The mother of Afonso II, the third king of Portugal, was Queen Urraca of Castile.
Correct Answer: Dulce Berenguer of Barcelona.
Returned by the tool: 1.—died March 25, 1223, Coimbra), the third king of Portugal (1211–23), under whom the reconquest of the south from the Muslims was continued. born: 1185?, Coimbra, Port. Afonso II was the son of King Sancho I and Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. 2…
Truth Guardian:FACTUAL. The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is correct. This is supported by the evidence that Afonso II was the third king of Portugal and that Queen Urraca of Castile was his mother. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Truth Seeker: NON-FACTUAL. The answer stating that the mother of Afonso II, the third king of Portugal, was Queen Urraca of Castile is incorrect. According to the provided knowledge, the mother of Afonso II was Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. Therefore, there is an incorrect conclusion in this question and answer.
Truth-Triangulator: As the Fact Verdict Manager, after conducting a thorough analysis of both perspectives and evidence, I have reached the following verdict:
NON-FACTUAL. The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is incorrect. The evidence provided shows that Afonso II was the son of King Sancho I and Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. Therefore, the answer contains false assumptions and incorrect conclusions. The Truth Seeker’s response is supported by the provided evidence, while the Proprietary Model’s response is not accurate as it did not consider the external evidence.
Query: Who is the Jeff Probst Show-nominated television producer’s first wife?
Response:His first wife was psychotherapist Shelley Wright.
Correct Answer: Shelley Wright.
Returned by the tool: 1.Probst was married to his first wife, psychotherapist Shelley Wright, from 1996 to 2001. In 2004, after filming on Survivor: Vanuatu wrapped, he began dating one of the contestants, a 23 year old Julie Berry - he was about 20 years her senior. They broke up in early 2008. 2. …
Truth Guardian: FACTUAL. The answer that Jeff Probst’s first wife was psychotherapist Shelley Wright is correct. According to the evidence provided, Jeff Probst was previously married to Shelley Wright, who is a *psychotherapist* in addition to being a television producer. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Truth Seeker: FACTUAL. According to the provided knowledge, Jeff Probst’s first wife was psychotherapist Shelley Wright. Therefore, the answer is correct and there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Truth-Triangulator: As the Fact Verdict Manager, after evaluating the responses from both AI models and examining the provided evidence, I have reached the following verdict:
FACTUAL. The answer that Jeff Probst’s first wife was psychotherapist Shelley Wright is correct. Both AI models correctly identified that there are no fallacies, faulty reasoning, or incorrect conclusions present in the question and answer. The evidence provided confirms that Jeff Probst was married to Shelley Wright from 1996 to 2001, and that she is a psychotherapist in addition to being a television producer. Therefore, the answer is factual and accurate.
Tool Enhancement
Considering the labor-intensive nature of tapping into external knowledge bases, we extend our exploration into exploiting ChatGPT’s potent contextual capabilities for tool-enhanced learning. Specifically inspired by preceding works such as (Chern et al., 2023; Qin et al., 2023), we harness the tool’s execution results to bolster large-scale models in hallucination detection. Employing a search engine as a tool to address queries from the “Query-Response” samples, we prompt ChatGPT to ascertain hallucination presence in the input “Query-Response” grounded on the search engine’s retrieved results, thereby providing concomitant justifications. Experimental outcomes, detailed in Table 3, not only showcase this method’s substantial improvement over the zero-shot approach predicated on ChatGPT but also further underscore the promising potential imbued in tool-enhanced learning.
Table 5. Ablation analysis on input context on FactCls.
Variants Conv. Multi. Comp. Set. Avg.
Alpaca-7B-LoRa 73.16 63.34 69.92 68.18 70.66
w/ facts 82.68 96.10 69.50 87.90 83.56
w/o query 39.0 10.04 9.24 9.52 30.48
4.3. A Strong Baseline: Triangulation for Truth
We categorize tool-enhanced ChatGPT as the Truth Seeker, which aims to make informed judgments by seeking external knowledge. However, it is inevitable that the information returned by external knowledge sources may be incomplete, erroneous, or redundant, thus potentially misleading the large-scale model. On the other hand, the detect-specific expert as the Truth Guardian relies on its own knowledge and expertise in the task, tending towards more conservative predictions. To address these challenges, we propose the Truth-Triangulator framework inspired by the “Triangulation for Truth” theory, involving verifying and confirming information by cross-referencing multiple independent sources or perspectives. Specifically, we fine-tune ChatGPT with over 10 examples as a Fact Verdict Manager and collect evidence from different angles or viewpoints to enhance the reliability and accuracy of the obtained truth or conclusion. Table 3 demonstrates that our model, compared to Lamma2-7b-chat-LoRA and GPT-3.5-turbo (tool), exhibits improvements. This finding further underscores the effectiveness of triangulation in mitigating errors and inconsistencies that may arise from relying on a single source or method, thereby facilitating a more comprehensive and robust understanding of the truth.
5. Analysis
5.0.1. Examining the Influences of Model Capacity.
As depicted in Figure 5, our insights reveal a notable enhancement in detecting fact-conflicting hallucinations when migrating from 7B to 13B models, especially within the realms of zero-shot and in-context learning scenarios. Notably, Alpaca-13B surpasses ChatGPT in performance, a phenomenon which can be ascribed to the implementation of consistent prompts throughout all experiments and a meticulous adjustment of prompts to cater to Alpaca’s heightened sensitivity in comparison to ChatGPT  [^3] Given that LLMs are acknowledged for their pronounced sensitivity to prompts, we do not assert that our prompts are universally optimal.. Intriguingly, when models are fine-tuned with training data, the impact of model capacity on performance amplification appears to be minimal. This observation intimates that training LLMs to function as hallucination detectors may effectively alleviate the necessity for augmented model capacities, thereby presenting an efficacious alternative pathway for enhancing model performance.
Figure 5. Impact of model size on hallucination detection.
5.0.2. Elevating Detection Capabilities: Analysis on Accurate and Rich Context
As discussed above, the incremental increase in model parameters for improving hallucination detection has shown diminishing returns, even up to the 13B model. To further investigate this, we orchestrate an ablation study aimed at scrutinizing the impact of implementing accurate facts for knowledge enhancement, as well as integrating query-based “Query-Response” context, to amplify the capabilities in hallucination detection. This investigation seeks to elucidate the nuanced role that precise and rich context plays in refining the model’s detection proficiency.
Enhancement lies in Accurate Fact.
We additionally investigate the straightforward application of the dataset’s intrinsic facts (i.e., the facts utilized during dataset construction) as retrieved evidence within the LoRA fine-tuning process. As illustrated in Table  5, incorporating factual information (w/ fact) culminates in a notable elevation in the FactCls score, substantiating the idea that the modest enhancement derived from retrieval is ascribed to the inferior quality of the garnered evidence. This also accentuates the considerable potential for augmentation by seeking the utmost accurate facts for evaluation.
Complete “Query-Response” Context.
A distinct characteristic of our dataset, which sets it apart from traditional fact-checking datasets, involves incorporating the “Query-Response” context. To delve deeper into its significance, we execute experiments, omitting the “query” during the Alpaca-7B-LoRA fine-tuning process. Table 5 displays the outcomes, wherein the “w/o query” scenario leads to a substantial decline of approximately 40% in the FactCls score, underscoring the pivotal role of a comprehensive “Query-Response” context in extracting invaluable information for informed decision-making.
5.0.3. Case Analysis.
To underscore the broad applicability of Truth-Triangulator, we embarked on supplementary tests, applying it to real-world hallucination data that transcends the boundaries of the FactCHD benchmark. We elucidate both the capabilities and constraints of our model by showcasing the findings from our out-of-distribution case analysis in Table 4. The results from these cases substantiate that our method can make proficient judgments, particularly when discrepancies arise between the detect-specific expert and tool-enhanced ChatGPT. This enhances the reliability of fact-conflicting hallucination detection by providing a practical examination of its application in genuine, unscripted scenarios.
6. Conclusion and Future Work
We introduce FactCHD, a meticulously designed benchmark tailored for the rigorous evaluation of fact-conflicting hallucinations within LLMs, which is notably enriched with a kaleidoscope of patterns and substantiated chain of evidence to fortify the robust elucidation of factuality assessments. Additionally, we unveil a semi-supervised data construction strategy, leveraging knowledge as facts to facilitate the creation of a hallucination dataset. Moreover, we delineate Truth-Triangulator, an innovative framework that elegantly employs the principle of triangulation to astutely discern the veracity of information, deploying cross-referencing generators and search tools to judiciously arbitrate responses, especially those teetering on the brink of questionable factuality.
Moving forward, our future endeavors will focus on three primary objectives: 1) explore efficient and powerful knowledge enhancement methods to enhance hallucination; 2) investigate factual errors stemming from incorrect or outdated prior knowledge, as well as the potential occurrence of hallucination during inference; 3) broadening our evaluative scope to encompass various modalities and cultural contexts in the assessment of hallucinations.
Appendix A In-depth Exploration into Benchmark Construction
Navigating through the space constraints, this section delicately provides a condensed, yet comprehensive overview of the data collection process, illustrated meticulously in Figure 6. We shed light on the intriguing prompt of ”Query-Response” generation, which seamlessly intertwines with four distinct actuality patterns, brought to life in Tables 6, 7, 8, and 9. Further, the meticulous design and specific prompt crucial for the chain of evidence generation are depicted in Table 10, offering a peek into the intricate mechanisms behind our methodology.
Appendix B A Glimpse into Experimental Intricacies
Pioneering through our experiment’s underpinnings, this section demystifies the prompts utilized for zero-shot and few-shot evaluation, articulated in Tables 11 and 12. Embarking further, we present an array of case examples extracted from the evaluative results of Alpaca-7B-LoRa, illustrated in Table13, along with a thoughtful elucidation of the palpable impact wielded by retrieval enhancement for Alpaca-7B-LoRA, specifically within the realm of Wikipedia exploration.
Figure 6. Knowledge as fact for constructing the benchmark on detecting fact-conflicting hallucination.
I want you to act as a question generator, your job is to generate questions with a *request* tone based on a given #Response#. Here are some examples:
#Response#: Chris Terrio is an American.
#Query#: Please tell me which country Chris Terri is from?
#Response#: The Renaissance was not a period in European history.
#Query#: Could you clarify whether the Renaissance was a period in European history?
#Response#: Quantum computing is a type of computation that uses quantum mechanics to process information.
#Query#: Would you kindly explain what quantum computing is?
You need to thoroughly study the above example to grasp the meaning of conventional Factuality. Please generate a complex and fluent query for the response without introducing any prior knowledge or the knowledge from the example. Make sure that the generated #Query# involve user instruction tone. Follow the example format for output:
#Response #: <Here is the claims with FACTUAL (SUPPORTS) or NON-FACTUAL (REFUTES) label from FEVER-like dataset and we need to leverage it for generating "Query-Response" contexts >
Table 6. Instruction for generating the “Query-Response” context based on FEVER-like dataset with the vanilla factuality pattern. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
I want you to act as a complex and fluent question and answer data generator, where your task is to generate complex and fluent questions and answers that require multi-hop reasoning based entirely on the given knowledge. Here are some examples:
#Knowledge#: ["Yao Ming", "spouse", "Ye Li"], ["Ye Li", "educated at", "Shanghai University of Sport"], ["Shanghai University of Sport", "establishment time", "November 1952"]
#Explanation#: To generate a question and answer, we need to perform multi-hop reasoning through the given knowledge. Here is the step-by-step reasoning:
1. Yao Ming’s wife is named Ye Li.
2. Ye Li was educated at Shanghai University of Sport.
3. Shanghai University of Sport was established in November 1952.
Based on the above reasoning chain, we can generate a complex question and answer with *3-hop* reasoning capabilities:
#Query#: Please tell me, when was the university that Yao Ming’s wife graduated from established?
#Correct response#: Yao Ming’s wife, Ye Li, graduated from Shanghai University of Sport, which was established in November 1952.
#Incorrect response#: Yao Ming’s wife, Ye Li, graduated from Shanghai University of Sport, which was established in March 1954.
<Demonstrations>
You need to thoroughly study the above example to grasp the meaning of multi-hop reasoning, the core is step-by-step reasoning*. Please generate a complex and fluent query and response that requires multi-hop reasoning based entirely on the given knowledge below without introducing any prior knowledge or the knowledge from the example. Make sure that the generated #Query# involve multi-hop reasoning and include *all* the information from the #Knowledge#! Response include #Correct response# and #Incorrect response#. Make ensure that the #Correct response# can be *correctly* deduced from the #Knowledge# using multi-hop reasoning without relying on unknown or insufficient information. #Incorrect response# sounds plausible but is factually incorrect. *To produce #Incorrect response#, you should choose one or more of the following strategies: fabricate information to resolve factual contradictions, misunderstand the question context and intention, provide an answer that is either too general or too specific, or employ incorrect reasoning to arrive at a hallucinated answer not supported by the knowledge. Follow the example format for output:
#Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
Table 7. Instruction for generating the “Query-Response” context based on KG with the multi-hop reasoning pattern. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
I want you to act as a complex and fluent question and answer data generator, where your task is to generate a complex and fluent question and answer that require quantitative comparison, relying solely on the provided knowledge. Here are some examples:
#Knowledge#: ["Chris Paul", "instance of", "human"], ["Chris Paul", "height", "183 centimetre"], ["Franklin Delano Roosevelt", "instance of", "human"], ["Franklin Delano Roosevelt", "height", "189 centimetre"]
#Query#: Could you kindly tell me who is taller, Chris Paul or Franklin Delano Roosevelt, both of whom are human?
#Correct response#: Franklin Delano Roosevelt is taller than Chris Paul. Roosevelt’s height is 189 centimeters while Chris Paul’s height is 183 centimeters.
#Incorrect response#: Chris Paul is taller than Franklin Delano Roosevelt. Chris Paul’s height is 183 centimeters while Roosevelt’s height is 189 centimeters.
<Demonstrations>
You need to thoroughly study the above example to grasp the meaning of quantitative comparison. Please generate a complex and fluent query and response that requires quantitative comparison based entirely on the given knowledge below without introducing any prior knowledge or the triplets from the example. Make sure that the generated #Query# involve quantitative comparison and include all the information from the #Knowledge#! Answers include #Correct response# and #Incorrect response#. Make ensure that the #Correct response# can be *correctly* deduced from the #Knowledge# using quantitative comparison without relying on unknown or insufficient information. #Incorrect response# sounds plausible but is factually incorrect. To produce #Incorrect response#, you should choose one or more of the following strategies: fabricate information to resolve factual contradictions, misunderstand the question context and intention, provide an answer that is either too general or too specific, or employ incorrect reasoning to arrive at a hallucinated answer not supported by the knowledge. Follow the example format for output:
#Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
Table 8. Instruction for generating the “Query-Response” context based on KG with the comparing fact. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
I want you to act as a complex and fluent question and answer data generator, where your task is to generate a complex and fluent question and answer that require set operation, relying solely on the provided knowledge. Here are some examples:
#Knowledge#: ["Mission: Impossible II", "producer", "Tom Cruise"], ["Mission: Impossible II", "original language of film or TV show", "English"], ["Mission: Impossible II", "composer", "Hans Zimmer"], ["The Last Samurai", "producer", "Tom Cruise"], ["The Last Samurai", "original language of film or TV show", "English"], ["The Last Samurai", "composer", "Hans Zimmer"]
#Query#:Could you provide the names of films produced by Tom Cruise, composed by Hans Zimmer, and with English as the original language of the film or TV show?
#Correct response#: The films that meet the given criteria are "Mission: Impossible II" and "The Last Samurai."
#Incorrect response#: The films that meet the given criteria are "Top Gun" and "Jerry Maguire."
<Demonstrations>
You need to thoroughly study the above example to grasp the meaning of set operation. Please generate a complex and fluent query and response that requires set operation based entirely on the given knowledge below without introducing any prior knowledge or the triplets from the example. Make sure that the generated #Query# involve set operation and include all the information from the #Knowledge#! Answers include #Correct response# and #Incorrect response#. Make ensure that the #Correct response# can be *correctly* deduced from the #Knowledge# using set operation without relying on unknown or insufficient information. #Incorrect response# sounds plausible but is factually incorrect. To produce #Incorrect response#, you should choose one or more of the following strategies: fabricate information to resolve factual contradictions, misunderstand the question context and intention, provide an answer that is either too general or too specific, or employ incorrect reasoning to arrive at a hallucinated answer not supported by the knowledge. Follow the example format for output:
#Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
Table 9. Instruction for generating the “Query-Response” context based on KG with the set operation. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
Table 10. Instruction to generate evidence-based chains of facts for the“Query-Response” contexts. The role description is indicated by the yellow text, while the factuality demonstration is represented by the purple text.
I want you to act as a data generator for a fallacy find task. You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in #Question# and #Answer#. I’ll tell you the right label: FACTUAL or NON-FACTUAL. Your job is to provide rational evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions that may be present in the #Question# and #Answer#. Here are some examples:
#Golden Label#: FACTUAL
#Query#: Can you tell me which country is Chris Terrios from?
#Response#: Chris Terrio is an American.
#Evidence#: [Chris_Terrio: Chris Terrio (born December 31, 1976) is an American screenwriter and film director.]
#Output#: FACTUAL. The answer that Chris Terrio is an American is correct. Chris Terrio is an American screenwriter and film director. So there are no fallacies, faulty reasoning, or incorrect conclusions present in this example.
#Golden Label#: NON-FACTUAL
#Query#: How many children did the head of government of the country where the distributor of "The Killer Inside Me" is located have?
#Response#: The head of government of the United States of America, John F. Kennedy, had 2 children.
#Evidence#: ["The Killer Inside Me", "distributor", "Warner Bros."], ["Warner Bros.", "owner of", "Warner Bros. Animation"], ["Warner Bros. Animation", "country", "United States of America"], ["United States of America", "head of government", "John F. Kennedy"], ["John F. Kennedy", "number of children", "4"]
#Output#: NON-FACTUAL. The answer stating that the head of government of the United States of America, John F. Kennedy, had 2 children is incorrect. "The Killer Inside Me" is distributed by Warner Bros., which is the owner of Warner Bros. Animation. Warner Bros. Animation is located in the United States of America, where the head of government was John F. Kennedy. However, John F. Kennedy had 4 children, not 2. Therefore, there is an incorrect conclusion in this query and response.
#Golden Lable#: <Here is the right label for the factuality of the following sample >
#Query#: <Here is the query for the LLM>
#Response#: <Here is the response from the LLM according to the above query>
#Evidence#: <Here is the evidence with fact to judge the factuality of above context>
#Output#:
I want you to act as a f̈allacy finder.̈ You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in the #Query# and the #Response#. Your job is to provide evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions which may be present in the #Query# and the #Response#. Begin your response with either FACTUAL or NON-FACTUAL, followed by your justification.
#Query #: <Here is the query for the LLM >
#Response #: <Here is the response from the LLM according to above query>
Table 11. Prompts for zero-shot evaluation. The role description is indicated by the yellow text.
I want you to act as a f̈allacy finder.̈ You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in the #Query# and the #Response#. Your job is to provide evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions which may be present in the #Query# and the #Response#. Begin your response with either FACTUAL or NON-FACTUAL, followed by your justification.
#Query#: Which city is London twinned with?
#Response#: London is twinned with *New York City*.
#Output#: FACTUAL.The response that London is twinned with New York City is correct. This is supported by the evidence that London is twinned with the administrative body of New York City. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this query and response.
#Response#: London is twinned with *New York City*.
#Query#: Which country has the highest total fertility rate in 1999?
#Response#: Zambia has the highest total fertility rate in 1999.
#Output#: The response that Zambia has the highest total fertility rate in 1999 is correct. According to the evidence provided, the total fertility rate for Zambia in 1999 was 6.116, which is the highest among all the countries listed. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this query and response.
#Query#: What is the IETF language tag for the variety of English used in L̈and of the Lost?̈
#Response#: The IETF language tag for the variety of English used in L̈and of the Lostïs *en-GB*.
#Output#: NON-FACTUAL.The response stating that the IETF language tag for the variety of English used in L̈and of the Lostïs en-GB is incorrect. L̈and of the Lostüses the English language, which has a part called Australian English. The IETF language tag for Australian English is en-AU, not en-GB. Therefore, there is an incorrect conclusion in this query and response.
#Query#: What is the IETF language tag for the part of English that Robert Young speaks, writes, or signs in?
#Response#: The IETF language tag for the part of English that Robert Young speaks, writes, or signs in is *en-GB*.
#Output#: NON-FACTUAL.The response stating that the IETF language tag for the part of English that Robert Young speaks, writes, or signs in is en-GB is incorrect. Robert Young speaks, writes, or signs in English, which has a part called Australian English. The IETF language tag for Australian English is en-AU, not en-GB. Therefore, there is an incorrect conclusion in this query and response.
# Query #: <Here is the query for the LLM>
# Response #: <Here is the response from the LLM according to above query>
Table 12. Prompts for few-shot evaluation. Note that the demonstration only adopted in ICL(4-shot) setting. Large language models are known to be highly sensitive to prompts, especially in zero-shot and in-context learning settings. Therefore, we cannot guarantee that the prompts we used in the evaluation of FACTLLM are optimal. The role description is indicated by the yellow text, the evaluation demonstration is represented by the purple text.
User Query Golden Label Prediction
Query: Could you tell me the name of the series that Rick Yune was on that got cancelled on December 12?
Response: Rick Yune was on a series that cancelled on December 12.
FACTUAL
FACTUAL. The answer that Rick Yune was on a series that got cancelled on December 12 is correct. Rick Yune was a cast member of the television series The Blacklist, which was cancelled on December 12, 2014. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Query: Could you inform me of the budget for the movie Poseidon?
Response:Poseidon had a budget of \$160 million.
NON-FACTUAL
NON-FACTUAL. The answer that Poseidon had a budget of \$160 million is incorrect. Poseidon had a *budget of \$150 million*, not \$160 million. Therefore, this is a false incorrect conclusion.
Query: Could you inform me of the budget for the movie Poseidon?
Retrieved: 1. Poseidon (film) - 7 million worldwide on a budget of \$160 million…… 2. Poseidon (DC Comics) - Poseidon is the name of a DC Comics deity who is based on the Greek god of the same name……
Response: Poseidon had a budget of \$160 million.
FACTUAL
FACTUAL. he answer that Poseidon had a budget of \$160 million is correct. The film had a budget of *160 million* and was released in the United States on May 12, 2006. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Table 13. Case example from the results of Alpaca-7B-LoRa (Tuning) and the illustration of the impact of retrieval enhancement.

### v2

@@
-Unveiling the Siren’s Song: Towards Reliable Fact-Conflicting Hallucination Detection
-Abstract.
-Large Language Models (LLMs), such as ChatGPT/GPT-4, have garnered widespread attention owing to their myriad of practical applications, yet their adoption has been constrained by issues of fact-conflicting hallucinations across web platforms.
-The assessment of factuality in text, produced by LLMs, remains inadequately explored, extending not only to the judgment of vanilla facts but also encompassing the evaluation of factual errors emerging in complex inferential tasks like multi-hop, and etc.
-In response, we introduce FactCHD, a fact-conflicting hallucination detection benchmark meticulously designed for LLMs.
-Functioning as a pivotal tool in evaluating factuality within “Query-Respons” contexts, our benchmark assimilates a large-scale dataset, encapsulating a broad spectrum of factuality patterns, such as vanilla, multi-hops, comparison, and set-operation patterns.
-A distinctive feature of our benchmark is its incorporation of fact-based chains of evidence, thereby facilitating comprehensive and conducive factual reasoning throughout the assessment process.
-We evaluate multiple LLMs, demonstrating the effectiveness of the benchmark and current methods fall short of faithfully detecting factual errors.
-Furthermore, we present Truth-Triangulator that synthesizes reflective considerations by tool-enhanced ChatGPT and LoRA-tuning based on Llama2, aiming to yield more credible detection through the amalgamation of predictive results and evidence [^1] The benchmark dataset and source code will be made available in https://github.com/zjunlp/FactCHD..
-Fact-Conflicting, Hallucination Detection, LLMs, Benchmark
-copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
-1. Introduction
-Figure 1. Illustrating Fact-Conflicting Hallucination Detection.
-There is a significant risk that LLMs (Zhao et al., 2023) may generate web content that appears credible but is factually incorrect or imprecise (Yin et al., 2023; Ji et al., 2023; Huang et al., 2023), thereby spreading misinformation on the web.
-This phenomenon is known as fact-conflicting hallucination (Zhang et al., 2023) and may result from flawed or outdated knowledge incorporated into the model’s parameters, as well as the language model’s shortcomings in tasks such as fact reconciliation, logical reasoning, and other sophisticated cognitive processes.
-This limitation not only restricts the application of generative AI in high-stakes areas such as finance, healthcare and law but also exacerbates the dissemination of erroneous information on the web community, posing a significant obstacle to the healthy development of the web.
-Given their pivotal role in ensuring the security of online content, effectively detecting and mitigating fact-conflicting hallucinations stands paramount.
-Traditional task definitions for fact verification (Wadden et al., 2020a, b) exhibit limitations, particularly in accommodating writing tasks that users frequently navigate when interfacing with LLMs.
-Moreover, existing hallucination evaluation benchmarks (Li et al., 2023a; Muhlgay et al., 2023), predominantly centering on rudimentary facts and textual content, lack in-depth exploration of complex operations among facts, thus rendering their coverage against fact-conflicting hallucinations suboptimal.
-In light of this, we introduce an inherently rigorous, yet authentic, task scenario: fact-conflicting hallucination detection, devoid of explicit claims or evidence.
-For instance, in Figure 1, when confronted with a query and its generated response, evaluators, within the prescribed scenario, are impelled to harness both their intrinsic knowledge and external resources, while employing reasoning skills to render a factual judgment accompanied by an elucidative reasoning process.
-In paving the way for future strides in hallucination evaluation, we dedicate our efforts towards establishing a rigorous benchmark, Fact-Conflicting Hallucination Detection (FactCHD), tailored for LLMs and encompassing a variegated array of factuality patterns, including Vanilla, Multi-hops, Comparison, and Set-Operation patterns.
-Drawing inspiration from the adage “to know it and to know the reason why of it” by Zhuzi, FactCHD endeavors to intertwine fact-based chains of evidence, providing cogent justifications when predicting the factual or non-factual nature of a statement.
-Achieving the intended goal of gathering a substantial amount of data through full-process human annotation is a time-consuming and resource-intensive task with limited scalability.
-Considering the high reliability of existing knowledge graphs (KGs) and textual knowledge, we advocate for a data construction strategy that leverages knowledge as facts, coupled with semi-supervised annotation, to facilitate the crafting of the aforementioned benchmark.
-As illustrated in Figure 3, we engineer our benchmark through a systematic construction strategy, bifurcated into two primary sections, encompassing six meticulous steps.
-Initially, a subset of queries is input into several LLMs, with corresponding responses being manually annotated, thereby curating a minimal but realistic “Query-Response” dataset enriched with factuality information.
-Subsequently, we intertwine knowledge sub-graphs and textual knowledge, utilizing these as factual underpinnings to guide ChatGPT in the autonomous generation of contextual “Query-Response” pairs.
-Then human similarity assessment is employed iteratively to finetune prompts, ensuring stylistic coherence between generated and realistic factuality data.
-The fourth step involves automated similarity screening of the synthesized factuality data, identifying and excising highly similar samples to uphold dataset diversity.
-Fifthly, ChatGPT is guided by precise instructions to generate evidence chains, anchoring its judgments in the “Query-Response” contexts on solid knowledge.
-Concludingly, an adept team of 21 individuals meticulously filters the data, adhering to stipulated rules, thereby assuring the procurement of a high-quality dataset.
-We conduct baseline experiments to assess the efficacy of several LLMs (such as Alpaca, ChatGPT, etc), utilizing our FactCHD under various settings—zero-shot, in-context learning, tuning for detect-specific expertise, and knowledge enhancement through retrieval/tools.
-Despite tuning and knowledge augmentation positively influencing the evaluation of fact-conflicting hallucinations, both open-source LLMs and ChatGPT encounter challenges in the precise and robust detection of factual inaccuracies.
-Thus, we introduce Truth-Triangulator, our innovative framework, inspired by the “Triangulation” theory (Valenza, 2016).
-Within this structure, tool-enhanced ChatGPT is designated as the Truth Seeker, and the detect-specific expert (based on Llama2-7B-LoRA) serves as the Truth Guardian.
-An additional agent, the Fact Verdict Manager, is tasked with amassing evidence from various perspectives to fortify the reliability and accuracy of the derived conclusions.
-Truth-Triangulator accentuates employing cross-referencing generators and search capabilities to astutely evaluate and adjudicate responses with potential factual discrepancies.
-Key insights are summarized as follows:
-  • We devise FactCHD, a comprehensive large-scale factuality evaluation benchmark that encompasses multiple domains and a myriad of factual dimensions, complemented by interpretable fact-based chains of evidence.
-This benchmark is poised to be an invaluable asset, propelling future innovations in detecting fact-conflicting hallucinations.
-  • We introduce a semi-supervised data construction strategy, utilizing knowledge as facts to expedite the assembly of the fact-conflicting hallucination dataset.
-Owing to the authenticity and extensive domain coverage of KGs, this strategy not only assures enhanced scalability but also stands out as particularly adept for large-scale applications.
-  • We present Truth-Triangulator, a framework employing triangulation for truth discernment, utilizing cross-referencing generators and search tools to arbitrate responses of dubious factuality.
-Initial experiments substantiate the varied performance of different LLMs in identifying fact-conflicting hallucinations and affirm the superiority of our proposed approach.
-2. Related Work
-Hallucination in LLMs.
-Despite the impressive capabilities of LLMs (Zhao et al., 2023; Chen et al., 2022b, a; Zhang et al., 2022), such as ChatGPT (OpenAI, 2022), in comprehending and carrying out user instructions, they sometimes generate misleading information with a high level of confidence.
-This phenomenon, known as hallucination (Ji et al., 2023; Ye et al., 2023), has been classified by (Zhang et al., 2023) into three categories: input-conflicting, context-conflicting, and fact-conflicting hallucinations.
-While humans can relatively easily identify the first two types, fact-conflicting hallucinations present a significant challenge as they involve the dissemination of misleading information on the web due to conflicting factual knowledge.
-While hallucination in vanilla natural language generation (NLG) settings has been extensively studied (Shuster et al., 2021; Creswell and Shanahan, 2022; Shuster et al., 2021; Das et al., 2023; Mallen et al., 2022), these methods primarily focus on analyzing hallucination across various NLP tasK.
-Li et al.
-(2023a) recently introduced HaluEval, the hallucination evaluation benchmark evaluating the performance of LLMs in recognizing hallucinations.
-Different from HaLuEval which only analyzes ChatGPT’s ability to evaluate hallucination, we specifically focus on the evaluation of fact-conflicting hallucination by constructing an interpretable benchmark that can serve as a public platform for checking the factual errors of context with explanation, as well as designing the Truth-Triangulator framework for reliable hallucination detection.
-Split #Sample Van. Multi. Comp. Set-Op.
-Train 51,383 31,986 8,209 5,691 5,497
-Test 6,960 4,451 1,013 706 790
-Table 1. Distribution of our proposed benchmark, FactCHD
-Table 2.
-Comparison with existing fact-checking datasets.
+FactCHD: Benchmarking Fact-Conflicting Hallucination Detection
+Abstract
+Despite their impressive generative capabilities, LLMs are hindered by fact-conflicting hallucinations in real-world applications.
+The accurate identification of hallucinations in texts generated by LLMs, especially in complex inferential scenarios, is a relatively unexplored area.
+To address this gap, we present FactCHD, a dedicated benchmark designed for the detection of fact-conflicting hallucinations from LLMs.
+FactCHD features a diverse dataset that spans various factuality patterns, including vanilla, multi-hop, comparison, and set operation.
+A distinctive element of FactCHD is its integration of fact-based evidence chains, significantly enhancing the depth of evaluating the detectors’ explanations.
+Experiments on different LLMs expose the shortcomings of current approaches in detecting factual errors accurately.
+Furthermore, we introduce Truth-Triangulator that synthesizes reflective considerations by tool-enhanced ChatGPT and LoRA-tuning based on Llama2, aiming to yield more credible detection through the amalgamation of predictive results and evidence [^1] The benchmark is available at https://github.com/zjunlp/FactCHD..
+1 Introduction
+Large Language Models (LLMs) Zhao et al.
+(2023) are susceptible to generating text that, while seemingly credible, can be factually inaccurate or vague, leading to the spread of misinformation online Yin et al.
+(2023); Ji et al.
+(2023); Huang et al.
+(2023).
+This issue, referred to as fact-conflicting hallucination Zhang et al.
+(2023), arises from the incorporation of incorrect or obsolete knowledge into the models’ parameters and from the models’ inherent limitations in complex cognitive ability.
+These shortcomings constrain LLMs’ deployment in critical domains like finance, healthcare, and law, and amplify the propagation of erroneous information.
+Figure 1: Illustration of fact-conflicting hallucination detection example from FactCHD, where the green part represents factual explanation core (body part) in the chain of evidence.
+Therefore, it is crucial to effectively detect fact-conflicting hallucinations for mitigating or editing them Yao et al. (2023).
+However, traditional fact verification tasks Wadden et al.
+(2020a, b) are not suitable for LLM-based “Query-Response” data.
+Moreover, existing hallucination evaluation benchmarks Li et al.
+(2023); Muhlgay et al.
+(2023), predominantly centering on vanilla facts and textual content, lack in-depth exploration of complex operations among facts, thus rendering their coverage against fact-conflicting hallucinations suboptimal.
+To bridge this gap, we introduce an inherently rigorous, yet authentic, task scenario: fact-conflicting hallucination detection, devoid of explicit claims or evidence.
+As shown in Figure 1, when confronted with a query and its generated response, detectors, are impelled to harness both their intrinsic knowledge and external resources, while rendering a factual judgment accompanied by an elucidative explanation.
+Datasets # Source Domains Construction Factuality Pattern Evaluation Expandability
+Fever Thorne et al. (2018) Text General HG Van. Consist.
+Climate-Fever Diggelmann et al. (2020) Text Climate Change HG Van. Consist.
+Health-Fever Sarrouti et al. (2021) Text Health HG Van. Consist.
+Sci-Fact Wadden et al. (2020a) Text Scientific HG Van. Consist.
+CoVERT Mohr et al. (2022) Text COVID-19 HG Van. Consist.
+TabFact Chen et al. (2020) Table General HG Van. Consist.
+HoVer Jiang et al. (2020) Text General HG Mul. Consist.
+FEVEROUS Aly and et al (2021) Text+Table General HG Van. Consist.
+HaluEval Li et al. (2023) Text General MIX Van. Hallucination
+FactCHD (ours) KGs (
+) & Text
+General&Vertical MIX Van.&Mul.&Com.&Set. Fact.+Chain.
+Table 1: Comparison with existing fact-checking datasets.
 Our FactCHD include both general and vertical domains, such as health, COVID-19, climate, science, and medicine (genes, virus and disease).
@@
 - chains of evidence.
-Datasets # Source Domains Construction Format Factuality Pattern Evaluation Expandability
-Fever (Thorne et al., 2018) Text General HG Claim Van. Consist.
-Climate-Fever (Diggelmann et al., 2020) Text Climate Change HG Claim Van. Consist.
-Health-Fever (Sarrouti et al., 2021) Text Health HG Claim Van. Consist.
-Sci-Fact (Wadden et al., 2020a) Text Scientific HG Claim Van. Consist.
-CoVERT (Mohr et al., 2022) Text COVID-19 HG Claim Van. Consist.
-TabFact (Chen et al., 2020) Table General HG Claim Van. Consist.
-HoVer (Jiang et al., 2020) Text General HG Claim Mul. Consist.
-FEVEROUS (Aly et al., 2021) Text+Table General HG Claim Van. Consist.
-HaluEval (Li et al., 2023a) Text General MIX Q&R Van. Hallucination
-FactCHD (ours) KGs (
-) & Text
-General&Vertical MIX Q&R Van.&Mul.&Com.&Set. Fact.+Chain.
+In paving the way for future strides in hallucination evaluation, we introduce a new benchmark, Fact-Conflicting Hallucination Detection (FactCHD), tailored for LLMs and encompassing a variegated array of factuality patterns, including Vanilla, Multi-hops, Comparison, and Set-Operation patterns.
+For example in Figure 1, querying whether the “Alps” or “Pyrenees” are higher, exemplifies a comparison pattern, assessing the relative relationships among facts.
+Drawing inspiration from the adage “to know it and to know the reason why of it” by Zhuzi, FactCHD extends beyond mere “Query-Response” labeling of hallucinations, incorporating golden chains of evidence to assess if detectors can provide coherent explanations for factualness judgment.
+Acknowledging the challenges of collecting comprehensive data through exhaustive human annotation, we propose a scalable data construction approach that harnesses existing knowledge graphs (KGs) and textual knowledge to create simulated hallucination instances with ChatGPT, verified with human annotation, for the efficient development of FactCHD.
+We evaluate the performance of various LLMs (such as Alpaca, Llama2-chat, and ChatGPT) using our FactCHD benchmark across multiple settings: zero-shot, in-context learning, specialized detection tuning, and knowledge enhancement via retrieval/tools.
+The results indicate that specialized detection tuning and knowledge enhancement notably improve the detection of fact-conflicting hallucinations.
+Additionally, we present Truth-Triangulator framework grounded in “Triangulation” theory Valenza (2016).
+This system comprises three roles: ChatGPT, enhanced with tools as the Truth Seeker; a detect-specific expert based on Llama2-7B-LoRA as the Truth Guardian; and the Fact Verdict Manager, which amasses evidence from another role to fortify the reliability and accuracy of the derived conclusions.
+Truth-Triangulator emphasizes the use of cross-referencing generators to astutely evaluate and adjudicate responses with potential factual discrepancies.
+Key insights are summarized as:
+  • We present FactCHD, a large-scale, multi-domain evaluation benchmark with diverse factual patterns and interpretable evidence chains, setting a new standard for detecting fact-conflicting hallucinations from LLMs.
+  • We introduce a scalable data construction strategy leveraging KGs, etc.
+to efficiently develop a fact-conflicting hallucination dataset, offering stronger applicability due to the authentic, broad domain coverage of KGs.
+  • We devise a triangulation-based framework Truth-Triangulator that employs cross-referencing generators for verifying LLM responses.
+Empirical results demonstrate the effectiveness of our approach in hallucination detection.
+2 Related Work
+Hallucination in LLMs.
+Figure 2: Overview of the factuality patterns invloved in our FactCHD.
+Figure 3: Domain distribution of FactCHD, where “General-p” and “General-d” denote common domains derived from Wikipedia and Wikidata, respectively.
+Despite LLMs Zhao et al.
+(2023); Chen et al.
+(2022); OpenAI (2022) like ChatGPT demonstrating remarkable understanding and execution of user instructions, they are prone to confidently generating misleading “hallucinations” Ji et al.
+(2023); Ye et al.
+(2023); et al (2023a); Hu et al.
+(2023).
+These hallucinations, as categorized by Zhang et al.
+(2023), can be input-conflicting, context-conflicting, or fact-conflicting—with the latter being especially problematic due to the propagation of inaccurate factual information online.
+While previous studies have extensively examined hallucinations within natural language generation (NLG) for various NLP tasks Shuster et al.
+(2021); Creswell and Shanahan (2022); Das et al.
+(2023); Mallen et al.
+(2022), HaluEval Li et al.
+(2023) has emerged as a recent benchmark for assessing LLMs’ recognition of such errors.
+Different from HaLuEval which only analyzes ChatGPT’s ability to evaluate whether hallucinatory, we specifically focus on the evaluation of fact-conflicting hallucination by constructing an interpretable benchmark that can serve as a public platform for checking the factual errors of context with explanation.
 Factuality Detection.
-Our work is also related to previous fact verification or actuality evaluation in various NLP tasks (Thorne et al., 2018; Wadden et al., 2020b; Gupta et al., 2022; Dziri et al., 2022; Rashkin et al., 2021; Kryściński et al., 2020; Pagnoni et al., 2021; Tang et al., 2022).
-Going beyond a simplistic binary notion of factuality, the FRANK framework (Pagnoni et al., 2021) advocates for a nuanced comprehension of factuality and introduces a typology of factual errors.
-Dhingra et al.
-(2019) measures $n$-gram lexical entailment for table-to-text generation.
-Recently,  Muhlgay et al.
-(2023) proposes FACTOR score describes LM’s tendency to assign greater likelihood to the original fact than to any of the non-factual alternatives.
-FELM (Chen et al., 2023) extends actuality evaluation into math and reasoning tasks.
-However, these benchmarks and metrics may not align well with the naturally occurring “Query-Response” format in the context of LLMs, and they emphasize accuracy evaluation while neglecting the provision of explanatory justifications for decision-making.
-In contrast, our work (1) provides challenge evaluation data in “Query-Response” format for LLMs; (2) considers an additional dimension of evaluation metrics that evaluate the interpretability of detecting fact-conflicting hallucination; and (3) investigates the effectiveness of Truth-Triangulator with verifying and confirming information by cross-referencing multiple independent sources or perspectives.
-Table 2 shows the comparison with other relevant datasets.
-Moreover, factuality-based hallucination detection establishes the foundation for model editing (Yao et al., 2023; Mao et al., 2023; Li et al., 2023b; Wang et al., 2023) of LLMs.
-3. FactCHD Benchmark Construction
-We first define the task of fact-conflicting hallucination detection and subsequently detail the benchmark construction process in the following subsections.
+Our research is also related to prior works on fact verification in NLP tasks Thorne et al.
+(2018); Wadden et al.
+(2020b); Gupta et al.
+(2022); Dziri et al.
+(2022); Liang et al.
+(2023); Rashkin et al.
+(2021); Kryściński et al.
+(2020), expanding the understanding of factuality beyond binary judgments.
+The FRANK framework Pagnoni et al.
+(2021) offers a detailed typology of factual errors, while  Dhingra et al.
+(2019) explores lexical entailment in text generation.
+The FACTOR score by Muhlgay et al.
+(2023) evaluates LMs based on the likelihood of factual content.
+Existing benchmarks often miss the “Query-Response” context of LLMs, focusing solely on accuracy without providing explanatory rationales.
+Our contribution lies in: (1) presenting “Query-Response” formatted data for LLMs; (2) introducing metrics for interpretability in detecting fact-conflicting hallucinations; and (3) validating the effectiveness of Truth-Triangulator through cross-referential verification from multiple sources.
+The comparison with other datasets is detailed in Table 1.
+3 Preliminaries
 Task Formulation.
-The detection of fact-conflicting hallucinations in LLM responses in real-world scenarios entails identifying factual inaccuracies and considering potential human-initiated requests.
-A detector should not only predict the factuality label accurately but also generate an explanatory justification for its binary prediction.
-Thus, the task of fact-conflicting hallucination detection is formally defined as: Input: A question $Q$ and response $R$ from LLMs, encompassing four types of fact-conflicting: Vanilla, Multi-hop, Comparison, and Set-Operation.
-Output: A sequence $A=[l,e]$, where $l$ and $e$ represent the factuality label (factual or non-factual) and explanation, respectively.
-Based on the aforementioned definitions, we construct FactCHD, which comprises a comprehensive array of training samples, enriched with an additional 6,960 meticulously curated samples for evaluating fact-conflicting hallucinations generated by LLMs.
-Our dataset ensures an equitable balance between factual or non-factual categories, providing a sturdy platform for evaluation.
-Notably, FactCHD presents three distinct characteristics: (1) It encompasses patterns of diverse reasoning, including multi-hop, comparison, and set operations as shown in Table 1; (2) FactCHD adheres to realistic scenarios by supplying “Query-Response” pairs and their associated evidence to validate the information provided; (3) The benchmark is meticulously crafted, leveraging KGs for initial data construction, followed by meticulous human verification to assure quality.
-Furthermore, the dataset inherently allows for expansion through updates based on KGs, offering a unique advantage in maintaining its contemporaneity and expandability.
-Further specific instructions and case data are illustrated in Appendix A.
-We proceed to introduce the design principles of the fact-conflicting hallucination detection benchmark in the following parts.
-Figure 2. Distribution of FactCHD across various domains, with “General-p” and “General-d” denoting prevalent domains sourced from Wikipedia and Wikidata, respectively.
-Addressing hallucinations in the outputs of LLMs necessitates a meticulous examination of both hallucinatory and non-hallucinatory instances.
-The susceptibility of LLMs to fact-conflicting hallucinations stems from their circumscribed knowledge and suboptimal reasoning capacities (Ji et al., 2023).
-Accordingly, we categorize four distinct patterns of factual errors, vividly illustrated in the lower section of Figure 3: (1) vanilla pattern that deals with factual statements that can be objectively verified using established sources and are typically more straightforwardly identified; (2) multi-hops pattern denotes the process of drawing conclusions by connecting multiple pieces of facts, (3) Comparison pattern refers to the act of evaluating and comparing the relative worth, and relationships between different pieces of facts, and (4) Set-operation pattern involves manipulating and combining sets of elements using operations to analyze relationships between different facts.
-Subsequently, we query open-source LLMs using specific instructions that reflect the identified factuality patterns.
-Through manual annotation of the hallucination responses, we curate a genuine compilation of hallucinations.
-Our objective is to gather a substantial amount of bulk hallucination data to simulate these realistic hallucinations as demonstrative examples.
-3.1. Generate “Query-Response” Contexts
-Capitalize knowledge as Fact
-KGs, with their extensive repositories of entities and relational data, not only champion compositional reasoning but also offer a foundational structure for factual information.
-Concurrently, textual knowledge, while invaluable for vanilla inference, takes on a pivotal role in offering additional and nuanced information.
-With this in mind, our aim steers toward exploring the capabilities of utilizing existent knowledge as a factual backbone in the semi-automated assembly of fact-conflicting hallucinations, detailed as follows:
-(1) We extract KG data from Wikidata (Vrandecic and Krötzsch, 2014) and PrimeKG (Chandak et al., 2023), utilizing it as an underpinning knowledge base for generating “Query-Respons” data.
-By manually selecting $438$ prevalent relations and crafting subgraphs through $K$-hop walks, performed $N$ times from divergent, randomly selected starting entities, we procure a varied set of subgraphs to be employed for multi-hop reasoning, fact comparison, and set operation.
-Heuristic rules are applied to ensure minimal intersection and coherence among the extracted subgraphs.
-(2) We adopt the text knowledge in these various fact verification datasets including FEVER (Thorne et al., 2018), Climate-Fever Diggelmann et al.
-(2020), Health-Fever (Sarrouti et al., 2021), COVID-FACT (Saakyan et al., 2021) and SCIFACT (Wadden et al., 2020a) to construct the data with vanilla pattern in our benchmark FactCHD.
-We only select the factual (supported) and non-factual (refuted) samples with their corresponding evidence and initially employ ChatGPT (GPT-3.5-turbo) to directly evaluate the claims in these datasets and select out samples where the model struggles to provide false answers.
-Figure 3. Overview of the construction and pattern illustration of FactCHD.
-3.2. Collect Realistic Data with Diverse Patterns
-Prompt for Q&R Generation
-We devise effective prompts to steer ChatGPT in generating the “Query-Response” contexts, comprising three vital elements: role description, factuality pattern, and factuality demonstration.
-The role description defines the system’s role and specifies the input and objective of the generation process.
-To ensure precise control over the type and quality of factuality-related samples, we give the explanation of corresponding factuality patterns and demonstrations to guide the model in generating “Query-Response” contexts.
-3.3. Refine Prompts with Consistency
-We utilize ChatGPT to generate “Query-Response” contexts, deploying manual prompts to enable its function as a role-playing agent, informed by specific demonstrative examples.
-Recognizing ChatGPT’s significant prompt sensitivity, we engage in an iterative refinement of them to ensure the generated data aligns closely with desired, realistic patterns, following the protocol delineated in § 3.1.
-Beginning with an initial fixed set of 100 examples, we select five samples for human similarity assessment, contrasting them with demonstrative examples.
-Employing the majority rule, we evaluate whether each context adheres to established similarity criteria.
-This iterative process allows for continuous prompt adjustment, assuring alignment between generated data patterns and desired aspirations.
-3.4. Automated Screening for Enhanced Diversity
-Aiming to improve the diversity of generated “Query-Response” contexts, we employ Sentence-BERT (SBERT)(Reimers and Gurevych, 2019) in an automated screening process, calculating average semantic similarity within contexts.
-This facilitates the identification and exclusion of highly similar samples, safeguarding dataset diversity.
-After the screening process, a total of 1,538 samples from the training set and 632 samples from the test set were removed, ensuring the final aggregate of samples is diverse.
-This meticulous elimination of semantically analogous entries ensures our collections exhibit varied and non-redundant queries and responses, enhancing the benchmark’s utility for evaluating diverse instances of fact-conflicting hallucination.
-3.5. Generate Chain of Evidence
-Our benchmark extends beyond merely identifying factual errors in “Query-Response” contexts produced by LLMs; it also necessitates the generation of coherent evidence chains to substantiate such judgments, all of which are rooted in factual knowledge and articulated by ChatGPT.
-Particularly, we exploit two types of knowledge sources (outlined in §3.1) as our factual foundation.
-ChatGPT employs either subgraph facts or textual facts to furnish comprehensive justifications when assigning appropriate labels to the “Query-Response” contexts.
-It’s pivotal to note that the efficacy of these explanations profoundly impacts the credibility and reliability of judgments, thereby fostering trust in predictive models.
-These explanatory outputs can also amplify users’ overall understanding.
-3.6. Quality Filtering via Human Annotator
-We devised filter rules encompassing three facets: pattern consistency, response factuality, and evidence chain logic, employing these to guide annotators through quality filtering.
-Our team of 21 annotators, each boasting proficient English reading skills and a minimum of undergraduate education, underwent uniform training per standardized annotation guidelines.
-The annotators leverage their own awareness and, when necessary, utilize search engines to ensure informed filtering decisions, discarding samples that flout established rules.
-Given the inherent subjectivity in defining matching quality via these rules, we divided both the annotators and samples into seven groups, each featuring three annotators reviewing the same data batch through a voting mechanism.
+Detecting fact-conflicting hallucinations in LLMs involves discerning factual errors in responses to human queries.
+A comprehensive detector must not only classify responses as factual or non-factual but also provide explanations for its judgments.
+We define the task as follows: Input: A question $Q$ paired with an LLM-generated response $R$, which may contain various fact conflicts.
+Output: A combined label and explanation sequence $A=[l,e]$, where $l$ is the binary factuality label (factual or non-factual), and $e$ articulates the rationale behind the assigned label.
+We evaluate the quality of $e$ using the $ExpMatch$ metric through the golden evidence chains in FactCHD.
+Factuality Patterns.
+We aim to explore distinct patterns of factual errors in our FactCHD, vividly illustrated in Figure 2.
+These include the (1) vanilla pattern dealing with factual statements that can be objectively verified using established sources, the (2) multi-hops pattern involving the process of concluding by connecting multiple pieces of facts, the (3) comparison pattern referring to the act of evaluating and comparing relative worth and relations between different pieces of facts, and the (4) set-operation pattern involving manipulating and combining sets of elements using operations to analyze relations between different facts.
+We generated corresponding “Query-Response” examples based on these patterns.
+Knowledge-Driven Factual Foundation for Reliable “Query-Response” Generation.
+KGs Liang et al.
+(2024) serve as a rich trove of structured entities and relations, ideal for compositional reasoning and anchoring factual data.
+Alongside this, textual knowledge is critical for nuanced inference beyond basic facts.
+Our research focuses on collecting existing knowledge and integrating it into prompts as a factual foundation for “Query-Respons” and chain of evidence generation, as outlined below: (1) We use $438$ widespread relations from Wikidata Vrandecic and Krötzsch (2014) and PrimeKG Chandak et al.
+(2023) to create varied subgraphs via $K$-hop walks, forming the knowledge base for generating “Query-Respons” examples with multi-hop reasoning, fact comparison, and set operation patterns.
+(2) We utilize text knowledge from datasets such as FEVER Thorne et al.
+(2018), Climate-Fever Diggelmann et al.
+(2020), Health-Fever Sarrouti et al.
+(2021), COVID-FACT Saakyan et al.
+(2021), and SCIFACT Wadden et al.
+(2020a) for generating examples with vanilla pattern.
+4 FactCHD Benchmark Construction
+Building on the aforementioned preliminaries, we develop FactCHD, a dataset containing a wealth of training instances and an additional 6,960 carefully selected samples for evaluating fact-conflicting hallucinations from LLMs.
+Our dataset maintains a balanced representation of factual and non-factual categories, offering a robust framework for assessment.
+The statistics and domain distribution of FactCHD are depicted in Table 2 and Figure 3.
+Next, we outline the design principles of our benchmark as follows.
+Split #Sample Van. Multi. Comp. Set-Op.
+Train 51,383 31,986 8,209 5,691 5,497
+Test 6,960 4,451 1,013 706 790
+Table 2: Data Statistic of our FactCHD
+4.1 Collect Realistic Data as Demonstration
+Subsequently, adhering to the defined factuality patterns, we manually craft corresponding queries and utilize open-source LLMs (such as ChatGLM THUDM (2023), Alpaca, and Vicuna) to generate authentic responses.
+By manually annotating these responses for hallucinations, we acquire “Query-Response” examples annotated with hallucination presence.
+Our objective is to curate a robust dataset of hallucination cases that closely emulate real-world examples, providing a demonstrative foundation and establishing standards for prompt refinement to ensure alignment between generated hallucinations and these demonstrations.
+4.2 Generate “Query-Response” Contexts
+Figure 4: Overview of the construction process of FactCHD.
+I want you to act as a data generator for a fallacy-finding task.
+You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in #Question# and #Answer#.
+I’ll tell you the right label: FACTUAL or NON-FACTUAL.
+Your job is to provide rational evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions that may be present in the #Question# and #Answer#.
+Here are some examples:
+#Golden Label#: NON-FACTUAL
+#Query#: How many children did the leader of the country, where the distributor of "The Killer Inside Me" operates, have?
+#Response#: The head of government of the United States of America, John F. Kennedy, had 2 children.
+#Evidence#: ["The Killer Inside Me", "distributor", "Warner Bros."], ["Warner Bros.", "owner of", "Warner Bros.
+Animation"], ["Warner Bros.
+Animation", "country", "United States of America"], ["United States of America", "head of government", "John F.
+Kennedy"], ["John F.
+Kennedy", "number of children", "4"]
+#Output#: NON-FACTUAL.
+The answer stating that the head of government of the United States of America, John F.
+Kennedy, had 2 children is incorrect.
+"The Killer Inside Me" is distributed by Warner Bros., which is the owner of Warner Bros.
+Animation.
+Warner Bros.
+Animation is located in the United States of America, where the head of government was John F.
+Kennedy.
+However, John F.
+Kennedy had 4 children, not 2.
+Therefore, there is an incorrect conclusion in this query and response.
+<Demonstrations>
+#Golden Lable#: <Here is the right label for the factuality of the following sample > #Query#: <Here is the query for the LLM> #Response#: <Here is the response from the LLM according to the above query> #Evidence#: <Here is the evidence with fact to judge the factuality of above context>
+#Output#:
+Table 3: Instruction to generate evidence-based chains of facts for the “Query-Response” contexts.
+The role description is indicated by the yellow text, while the factuality demonstration is represented by the purple text.
+Generate “Query-Response” based on Knowledge with ChatGPT.
+We incorporate the knowledge collected on distinct factuality patterns into customized prompts, specifying whether to generate factual or non-factual responses.
+This guides ChatGPT in producing “Query-Respons” instances across various factuality categories with golden factuality labels.
+Detailed prompts are in Table 7 through  10 of  §A.
+Refine Prompts with Consistency.
+In the initial phase of generating “Query-Response” instances, we assess the consistency of five samples per pattern against demonstrations, using majority consensus.
+We subjectively evaluate each context’s adherence to the style and form of realistic demonstrations, employing iterative refinements of prompts to ensure that the generated data exhibits realistic patterns.
+We aim for an at least 95% consistency rate before scaling up the generation of “Query-Response” instances.
+4.3 Automated Semantic Diversity Screening
+To increase “Query-Response” context diversity, we employ Sentence-BERT (SBERT)Reimers and Gurevych (2019) to automatically compute semantic similarity matrices, filtering out near-duplicate samples to preserve dataset variety.
+During filtration, we removed 1,542 training and 832 test samples, guaranteeing a diverse final dataset.
+This careful pruning of semantically similar entries promotes a varied collection of queries and responses, bolstering the benchmark’s effectiveness for assessing diverse instances.
+4.4 Generate Chain of Evidence
+Our benchmark evaluates the detectors’ ability to not only identify hallucinations but also to provide effective explanations.
+It necessitates ChatGPT’s generation of coherent golden evidence chains grounded in factual knowledge for substantiating judgments.
+Utilizing subgraph or textual facts outlined in §section 3, along with the previously generated “Query-Response” pairs, ChatGPT delivers thorough justifications for labels assigned to “Query-Response” contexts.
+Detailed prompts are presented in Tables 3 These golden evidence chains are critical for assessing the explanatory validity of hallucination detectors.
+4.5 Quality Filtering via Human Annotator
+We craft filter rules to ensure pattern consistency, response factuality, and logical evidence chains for quality control in the annotation.
+21 educated annotators are uniformly trained and utilize both their expertise and search tools for rigorous sample vetting.
+To minimize subjectivity, we divided them into seven triads, introducing a voting mechanism for data evaluation.
 Simultaneously judged mismatches by annotators led to sample discarding, resulting in final removal counts of 565 and 258 samples from the training and test sets, respectively.
-4. Experiments
-[…]
+We involve Fleiss’s Kappa ($\kappa$) as a measure of inter-annotator agreement to assess the reliability of our annotations.
+We calculate $\kappa$ over a set of 100 annotated samples, resulting in a $\kappa$ value of 0.808, indicating substantial agreement.
 Evaluator Vanilla Multi-hops Comparison Set-Operation Average
@@
 ICL (4-shot)
+GPT-3.5-turbo 62.02 37.29 65.66 51.85 32.2 48.11 64.74 50.14 $\uparrow$8.2261.04 $\uparrow$17.9341.96
+text-davinci-003 56.52 39.36 55.02 58.22 8.50 48.53 50.34 51.82 $\uparrow$1.952.88 $\uparrow$24.8844.45
+Alpaca-7B 35.82 31.01 18.12 40.16 8.86 29.28 6.70 31.52 $\uparrow$5.2428.34 $\uparrow$16.7632.23
+Vicuna-7B 41.36 42.51 29.24 58.35 19.36 41.55 13.46 53.60 $\uparrow$6.335.14 $\uparrow$19.1445.98
+Llama2-7B-chat 31.00 39.08 39.13 54.38 10.50 41.83 27.96 51.73 $\uparrow$24.4830.25 $\uparrow$13.6143.02
+Det.(tune)
+Alpaca-7B-LoRA 73.14 49.00 63.34 70.83 69.92 59.88 68.18 63.75 $\uparrow$42.3270.66 $\uparrow$22.7354.96
+Vicuna-7B-LoRA 73.52 48.07 64.72 71.74 67.34 62.08 50.36 66.04 $\uparrow$34.4469.58 $\uparrow$9.0054.98
+Llama2-7B-chat-LoRA 77.41 47.91 67.70 67.30 62.27 57.03 78.68 65.94 $\uparrow$44.4874.73 $\uparrow$10.6953.71
+Knowledge
+Alpaca-7B-LoRA (wiki) 73.86 49.44 67.3 69.97 68.24 60.25 67.38 63.00 $\uparrow$0.6671.32 $\uparrow$0.1155.07
+Vicuna-7B-LoRA (wiki) 75.14 49.56 65.46 72.71 65.10 63.51 55.42 66.65 $\uparrow$1.28 70.86 $\uparrow$1.3056.28
+Llama2-7B-chat-LoRA (wiki) 77.14 46.71 69.61 64.17 66.05 49.73 78.08 64.52 $\uparrow$1.13 75.86 $\uparrow$0.8754.58
+GPT-3.5-turbo (tool) 69.71 38.60 69.92 48.43 44.08 47.26 74.21 45.65 $\uparrow$7.5968.63 $\uparrow$0.4142.37
+Truth-Triangulator 80.97 47.08 75.01 64.21 66.27 55.70 80.87 65.25 78.15 52.52
 […]
-[…]
-[…]
-[…]
-[…]
-Det.(tune)
-[…]
-[…]
-[…]
-Knowledge
-[…]
-[…]
-[…]
-[…]
-Truth-Triangulator 80.97 47.08 75.01 64.21 66.27 55.70 80.87 65.25 78.15 52.52
-4.1. Metrics and Settings
-FactCls Metric
-We employ the FactCls metric, denoted by the Micro F1 score, to assess binary factuality classification performance, with a distinct focus on the distribution $p(l|Q\&R)$, classifying instances as either factual or non-factual.
-In light of our commitment to pinpointing non-factual examples, we designate non-factual as the positive class and factual as the negative class.
-ExpMatch Metric
-In FactCHD, the chain of evidence integrates two elements: introductory, expository statements and a core of factual explanations.
-The former provides phrasing like ’Therefore, there is an incorrect conclusion in this query and response’ as a segue, while the latter delivers detailed insights, bolstering the reasoning behind factuality judgments.
-Given the paramount importance of aligning factual explanations over expository parts, we employ a segmented matching and weighted averaging metric, ExpMatch, amalgamating these sections for evaluative precision.
-Specifically, we determine $\text{Score}_{bd}$ using span-based Micro F1 to measure word overlap at the unigram level between the model-generated and reference bodies, reflecting factual core matching.
-Concurrently, $\text{Score}_{ht}$ is defined via the ROUGE-L score for the head-to-tail part, assessing similarity based on the longest common word subsequence between system-generated context and gold reference.
-Ultimately, we introduce the ExpMatch metric, a composite of $\text{Score}_{bd}$ and $\text{Score}_{ht}$, serving as the pivotal metric for appraising the coherence and quality of utilized factual chains in explanations:
-(1) $\textsc{ExpMatch}=\alpha\times\text{Score}_{bd}+(1-\alpha)\times\text{Score}_{ht}.$
-In our experiment, we initially assign $\alpha$ to 0.7.
-If a generated result lacks factual or non-factual information, indicating a prediction failure, we attribute an ExpMatch value of 0 to that particular example during metric calculation.
-4.2. Empirical Examination of Baseline Strategies
-This study seeks to probe the efficiency and viability of several prototypical baseline strategies in the detection of fact-conflicting hallucinations.
-These strategies encompass: (1) Zero-shot Learning, which assesses model performance without prior training in hallucination detection; (2) In-context Learning, employing 4-shot samples as demonstrations to prompt the model and evaluate its competence in managing context-specific hallucinations; (3) Detect-Specific Expert Model, finessing the LoRA (Hu et al., 2022) parameters through fine-tuning, utilizing our specialized training set to harness domain expertise; (4) Knowledge Enhancement, amplifying the model’s capabilities by integrating external knowledge via retrieval techniques or tool-based enhancements.
-Through these experiments, we aim to traverse both the potentials and limitations of these baseline methods within the scope of hallucination detection.
+5 Experiments
+5.1 Metric Definition
+FactCls Metric.
+We employ the FactCls, denoted by the Micro F1 score, to evaluate binary factuality classification performance.
+This metric focuses on the distribution $p(l|Q\&R)$, classifying instances as either factual or non-factual.
+With a specific emphasis on identifying non-factual examples, we designate non-factual as the positive class and factual as the negative class.
+ExpMatch Metric.
+In FactCHD, the golden evidence chain features introductory/expository statements (head-tail part) and a factual explanation core (body part).
+The former contextualizes with phrases such as ‘Therefore, there is an incorrect conclusion in this query and response’, and the latter delivers the in-depth, fact-based reasoning process.
+For hallucination detectors, prompts guide their outputs akin to the gold evidence chain for quality assessment.
+Given the paramount importance of aligning factual explanations over expository parts, we introduce ExpMatch, a metric employing segmented matching with weighted averaging.
+It computes $\text{Score}_{bd}$ via span-based Micro F1 for unigram overlap between generated and reference bodies, and $\text{Score}_{ht}$ via ROUGE-L to assess similarity based on the longest common word subsequence for the head-to-tail part.
+ExpMatch, combining $\text{Score}_{bd}$ and $\text{Score}_{ht}$, evaluates the explanations generated from detectors as:
+$\small\textsc{ExpMatch}=\alpha\times\text{Score}_{bd}+(1-\alpha)\times\text{Score}_{ht}.$ (1)
+We initially set $\alpha$ to 0.7, emphasizing the body part [^2] Preliminary evaluations reveal that scores with $\alpha$ at 0.75 and 0.8 align with our established conclusions.
+If a generated result lacks factual or non-factual information, indicating a prediction failure, we attribute an ExpMatch value of 0 to that particular example during metric calculation..
+5.2 Experimental Settings
 Evaluation Models.
-We evaluate the performance of various prominent LLMs within the context of our FactCHD benchmark, primarily focusing on models accessible through the OpenAI API, such as text-davinci-003, which corresponds to the InstructGPT model trained using supervised instruction tuning (Ouyang et al., 2022), as well as GPT-3.5-turbo (ChatGPT), which represents one of the most widely recognized models.
-Alternatively, we explore the adoption of open-source models such as Llama2-chat (Touvron et al., 2023b), Alpaca (Taori et al., 2023) and Vicuna (Chiang et al., 2023), which are fine-tuned variants of the LLaMA (Touvron et al., 2023a).
-Implementation Details.
-We conduct the generation of factuality evaluation instances using the Azure OpenAI ChatGPT API [^2] https://openai.com/blog/chatgpt/.
-During the generation process, we employ a temperature of $1.0$ to control the diversity of generated samples, while limiting the maximum number of tokens to $2048$ to ensure concise responses.
-Additionally, we set the frequency penalty to zero and utilize a Top-$p$ value of $1.0$ to ensure unrestricted token selection during generation.
-For evaluation purposes, we set the temperature to 0.2 across all models to reduce output randomness.
-Specific instructions employed in experiments as detailed in Appendix B.
-Figure 4.
-Overview Truth-Triangulator.
+We evaluate various leading LLMs on FactCHD benchmark, focusing on OpenAI API models, including text-davinci-003 (InstructGPT) and GPT-3.5-turbo (ChatGPT).
+Additionally, we explore the adoption of open-source models such as Llama2-chat et al (2023b), Alpaca Taori et al.
+(2023) and Vicuna Chiang et al.
+(2023), which are fine-tuned variants of the LLaMA Touvron et al.
+(2023).
+Figure 5: Overview Truth-Triangulator.
 Here we designate the Refer to caption as the “Truth Guardian” based on Llama2-7B-chat-LoRA while Refer to caption as the “Truth Seeker” based on GPT-3.5-turbo (tool) in our experiments.
 We want the Refer to caption “Fact Verdict Manager” to collect evidence from different viewpoints to enhance the reliability and accuracy of the obtained conclusion.
-4.2.1. Zero-shot Learning Performance
-The empirical results of the evaluation, presented in Table 3, reveal that LLMs with zero-shot learning still struggle with identifying implicit factuality in “Query-Response” contexts.
-The performance of the open-source 7B LLMs in zero-shot learning is notably poor, highlighting its deficiency in both instruction comprehension and internal knowledge representation.
-Even the ChatGPT exhibits limited proficiency in distinguishing between factual and non-factual “Query-Response” samples, achieving a mere 52.82% FactCls score in zero-shot, just slightly above chance level.
-Further investigation into the underwhelming performance reveals a notable drawback: LLMs (especially Llama2-chat) are susceptible to false negatives, resulting in reduced sensitivity when it comes to error detection.
-4.2.2. In-context Learning Performance
-[…]
-4.2.3. Detect-Specific Expert Performance
-This study explores the efficacy of tuning LLMs using LoRA to serve as domain-specific experts in identifying fact-conflicting hallucinations.
-Through training LLMs with our dedicated instructions and datasets, which are intrinsically focused on factuality, we aim to hone their proficiency in this particular domain.
-Our results highlight a notable enhancement across all open-source 7B models, wherein Llama2-chat emerges as the top-performing model amongst its peers.
-Post-fine-tuning with fact-conflicting hallucination data, Llama2-chat-7B models attain peak performance on our benchmark, as evidenced by a 74.73% FactCls score and a 53.71% ExpMatch score, respectively.
-The demonstrated success in training a scaled-down 7B model using our approach underlines the potential of employing 7B LLMs as effective hallucination detectors and paves the way for additional research and trials in this field.
-4.2.4. Knowledge Enhancemant
-Retrieval Enhancement
-Leveraging retrieval of pertinent knowledge stands as a prevalent strategy to mitigate factual inaccuracies (Lewis et al., 2020).
-This study empowers LLMs with knowledge facts derived from Wikipedia, deploying TF-IDF for document retrieval and subsequently applying BM25 to pinpoint the top two most pertinent paragraphs from the fetched document.
-Given the 7B model’s constrained capacity to comprehend lengthy instructional texts—as unveiled in our ICL experiments—we elevate its grasp of expansive knowledge by supplying knowledge for fine-tuning the LoRA parameters.
+Implementation Details.
+Using Azure’s OpenAI ChatGPT API [^3] https://openai.com/blog/chatgpt/, we generate samples with a temperature of $1.0$ to control the diversity of generated samples, while limiting the maximum number of tokens to $2048$ to ensure concise responses.
+We use a frequency penalty of zero and a Top-$p$ of $1.0$ to ensure unrestricted token selection during generation.
+For evaluations, we standardize the temperature at $0.2$ to minimize randomness.
+Detailed instructions are provided in Table 11 through Table 12 of §B.
+Baseline Strategy Settings
+This study investigates the effectiveness of various baseline strategies in detecting fact-conflicting hallucinations.
+These strategies include: (1) Zero-shot Learning, which evaluates model performance without prior training in hallucination detection; (2) In-context Learning, using 4-shot samples as demonstrations to prompt the model and evaluate its ability to handle hallucination detection; (3) Detect-Specific Expert Model, fine-tuning the LoRA Hu et al.
+(2022) parameters using our specialized training set to leverage domain expertise; (4) Knowledge Enhancement, enhancing the model’s capabilities by integrating external knowledge through retrieval techniques or tool-based enhancements.
+These experiments aim to explore both the potential and limitations of these baseline methods in hallucination detection.
+5.3 Empirical Experiment Results
+5.3.1 Zero-shot Learning Performance
+The empirical results presented in Table 4 show that LLMs with zero-shot learning struggle to identify implicit factuality in “Query-Response” contexts.
+The performance of the open-source 7B LLMs in zero-shot learning is notably poor, indicating deficiencies in both instruction comprehension and internal knowledge representation.
+Even the ChatGPT shows limited proficiency in distinguishing between factual and non-factual “Query-Response” samples, achieving only a 52.82% FactCls score in zero-shot.
+Llama2-chat is susceptible to false negatives, resulting in reduced sensitivity in hallucination detection.
+5.3.2 In-context Learning Performance
+The incorporation of few-shot information significantly improves fact-conflicting hallucination detection in the GPT-3.5-turbo, Alpaca-7B, Vicuna-7B, and Llama2-7B-chat models.
+Compared to models operating without this additional context, these enhancements result in an average increase of approximately $\uparrow$6in the FactCls and $\uparrow$18in ExpMatch scores.
+However, the impact of integrating few-shot information on text-davinci-003 is relatively modest, indicating its limited proficiency in managing few-shot learning.
+Furthermore, the improvement from incorporating few-shot information into Llama2-chat is significant, possibly because of the model’s superior contextual learning abilities compared to Alpaca and Vicuna, leading to a better understanding of few-shot demonstrations.
+5.3.3 Detect-Specific Expert Performance
+We investigate the effectiveness of tuning LLMs with LoRA for domain-specific expertise in detecting fact-conflicting hallucinations.
+By training these models on our trainset, we enhance their detection task proficiency.
+Our findings show that all tested open-source 7B models improve, with Llama2-chat outperforming its counterparts.
+After fine-tuning on hallucination data, Llama2-chat-7B achieves a FactCls score of 74.73% and an ExpMatch score of 53.71% on our benchmark.
+This success with a 7B model underscores the viability of using such models as hallucination detectors and encourages further exploration in this area.
+5.3.4 Knowledge Enhancemant
+Retrieval Enhancement.
+This study enhances LLM-based detection with Wikipedia-sourced facts, using BM25 for initial document retrieval to select the top five most relevant paragraphs.
+Our experiments show that given the 7B model’s limited ability to process long texts, providing knowledge during fine-tuning modestly improves performance.
 However, compared to the detect-specific expert without knowledge augmentation, the improvement achieved by combining Wikipedia retrieval with fine-tuning the LoRA parameters is relatively modest.
-We attribute these observations to two primary factors: our dataset spans multiple domains while Wikipedia knowledge encompasses merely a subset, and our retrieval method, being elementary and generic, occasionally yields lower-quality evidence.
-Table 4. Out-of-distribution case analysis of our Truth-Triangulator.
-User Contexts and Returned Evidence Detection Opinion from Both Sides Prediction with Triangulation
+We attribute these observations to two primary factors: our dataset spans multiple domains while Wikipedia encompasses merely a subset, and our retrieval method being elementary, occasionally yields lower-quality evidence.
+Query and Returned Evidence Detection Opinion from Both Sides Prediction with Triangulation
 Query: Who was the mother of Afonso II, the third king of Portugal?
@@
 Therefore, there is an incorrect conclusion in this question and answer.
-Truth-Triangulator: As the Fact Verdict Manager, after conducting a thorough analysis of both perspectives and evidence, I have reached the following verdict:
-NON-FACTUAL.
+Truth-Triangulator: As the Fact Verdict Manager, after conducting a thorough analysis of both perspectives and evidence, I have reached the following verdict: NON-FACTUAL.
 The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is incorrect.
@@
 Correct Answer: Shelley Wright.
-Returned by the tool: 1.Probst was married to his first wife, psychotherapist Shelley Wright, from 1996 to 2001.
+Returned by the tool: 1.
+Probst was married to his first wife, psychotherapist Shelley Wright, from 1996 to 2001.
 In 2004, after filming on Survivor: Vanuatu wrapped, he began dating one of the contestants, a 23 year old Julie Berry - he was about 20 years her senior.
@@
 Therefore, the answer is factual and accurate.
-Tool Enhancement
-Considering the labor-intensive nature of tapping into external knowledge bases, we extend our exploration into exploiting ChatGPT’s potent contextual capabilities for tool-enhanced learning.
-Specifically inspired by preceding works such as (Chern et al., 2023; Qin et al., 2023), we harness the tool’s execution results to bolster large-scale models in hallucination detection.
-Employing a search engine as a tool to address queries from the “Query-Response” samples, we prompt ChatGPT to ascertain hallucination presence in the input “Query-Response” grounded on the search engine’s retrieved results, thereby providing concomitant justifications.
-Experimental outcomes, detailed in Table 3, not only showcase this method’s substantial improvement over the zero-shot approach predicated on ChatGPT but also further underscore the promising potential imbued in tool-enhanced learning.
-Table 5. Ablation analysis on input context on FactCls.
+Table 5: Case analysis of out-of-distribution examples from ChatGPT using Truth-Triangulator.
+Tool Enhancement.
+Considering the labor-intensive nature of tapping into external knowledge bases, we investigate leveraging ChatGPT’s advanced contextual understanding for tool-enhanced hallucination detection, drawing from prior research such as Chern et al.
+(2023); Qin et al.
+(2023).
+By using a search engine to inform ChatGPT’s analysis of “Query-Response” samples, we enable the model to identify hallucinations with supporting evidence from search results.
+Results in Table 4 demonstrate significant enhancements over ChatGPT’s few-shot performance, highlighting the efficacy of tool learning.
+5.4 Exploring Triangulation for Truth
+We categorize tool-enhanced ChatGPT as the Truth Seeker, which aims to make informed judgments by seeking external knowledge.
+However, the information returned by external knowledge sources may inevitably be incomplete, erroneous, or redundant, thus potentially misleading the large-scale model.
+On the other hand, the detect-specific expert as the Truth Guardian relies on its knowledge and expertise in the task, tending towards more conservative predictions.
+To address these challenges, we propose the Truth-Triangulator framework inspired by the “Triangulation for Truth” theory, involving verifying and confirming information by cross-referencing multiple independent perspectives.
+Figure 5 illustrates our approach of fine-tuning ChatGPT as a Fact Verdict Manager, leveraging the output and knowledge from Truth Seeker and Truth Guardian to boost conclusion reliability.
+Table 4 reveals our model’s superior performance to Lamma2-7b-chat-LoRA and GPT-3.5-turbo (tool), highlighting the benefits of triangulation in reducing single-source error and enhancing truth verification.
+5.5 Experimental Analysis
+Examining the Influences of Model Capacity.
+Figure 6 illustrates that transitioning from 7B to 13B models notably improves the detection of fact-conflicting hallucinations, particularly in zero-shot and in-context learning scenarios.
+Alpaca-13B outperforms ChatGPT, which can beascribed to the consistently adopted command prompt may be more friendly to Alpaca-13B  [^4] We keep consistent prompts for each LLM throughout all experiments.
+Given that LLMs are acknowledged for their pronounced sensitivity to prompts, we do not assert that our prompts are universally optimal..
+Interestingly, when models are fine-tuned with training data, the impact of model capacity on performance improvement appears minimal.
+This implies that further training larger LLMs as hallucination detectors may yield limited benefits, and it is necessary to explore alternatives with higher upper limits for enhancing detection performance.
+Figure 6: Analysis of model capacity impact.
+Enhancement lies in Accurate Evidence.
+We also explore the direct use of intrinsic facts from the dataset as golden evidence in the LoRA fine-tuning process.
+As shown in Table  6, integrating factual information (w/ gold evidence) leads to a significant improvement in the FactCls score, indicating that the modest improvement from retrieval may stem from the lower quality of the obtained evidence.
+This underscores the potential for substantial improvement by seeking the most accurate facts for evaluation.
+Complete “Query-Response” Context.
+Our dataset differs from traditional fact-checking datasets in its inclusion of the “Query-Response” context.
+To examine its impact, we conduct experiments by excluding the “query” during the Alpaca-7B-LoRA fine-tuning process.
+The results in Table 6 reveal that the “w/o query” scenario leads to a significant 40% decrease in the FactCls score, highlighting the essential role of a comprehensive “Query-Response” context in extracting valuable information for informed decision-making.
 Variants Conv. Multi. Comp. Set. Avg.
 Alpaca-7B-LoRa 73.16 63.34 69.92 68.18 70.66
-w/ facts 82.68 96.10 69.50 87.90 83.56
+w/ golden evidence 82.68 96.10 70.50 87.90 83.56
 w/o query 39.0 10.04 9.24 9.52 30.48
-4.3. A Strong Baseline: Triangulation for Truth
-We categorize tool-enhanced ChatGPT as the Truth Seeker, which aims to make informed judgments by seeking external knowledge.
-However, it is inevitable that the information returned by external knowledge sources may be incomplete, erroneous, or redundant, thus potentially misleading the large-scale model.
-On the other hand, the detect-specific expert as the Truth Guardian relies on its own knowledge and expertise in the task, tending towards more conservative predictions.
-To address these challenges, we propose the Truth-Triangulator framework inspired by the “Triangulation for Truth” theory, involving verifying and confirming information by cross-referencing multiple independent sources or perspectives.
-Specifically, we fine-tune ChatGPT with over 10 examples as a Fact Verdict Manager and collect evidence from different angles or viewpoints to enhance the reliability and accuracy of the obtained truth or conclusion.
-Table 3 demonstrates that our model, compared to Lamma2-7b-chat-LoRA and GPT-3.5-turbo (tool), exhibits improvements.
-This finding further underscores the effectiveness of triangulation in mitigating errors and inconsistencies that may arise from relying on a single source or method, thereby facilitating a more comprehensive and robust understanding of the truth.
-5. Analysis
-5.0.1. Examining the Influences of Model Capacity.
-As depicted in Figure 5, our insights reveal a notable enhancement in detecting fact-conflicting hallucinations when migrating from 7B to 13B models, especially within the realms of zero-shot and in-context learning scenarios.
-Notably, Alpaca-13B surpasses ChatGPT in performance, a phenomenon which can be ascribed to the implementation of consistent prompts throughout all experiments and a meticulous adjustment of prompts to cater to Alpaca’s heightened sensitivity in comparison to ChatGPT  [^3] Given that LLMs are acknowledged for their pronounced sensitivity to prompts, we do not assert that our prompts are universally optimal..
-Intriguingly, when models are fine-tuned with training data, the impact of model capacity on performance amplification appears to be minimal.
-This observation intimates that training LLMs to function as hallucination detectors may effectively alleviate the necessity for augmented model capacities, thereby presenting an efficacious alternative pathway for enhancing model performance.
-Figure 5. Impact of model size on hallucination detection.
-5.0.2. Elevating Detection Capabilities: Analysis on Accurate and Rich Context
-As discussed above, the incremental increase in model parameters for improving hallucination detection has shown diminishing returns, even up to the 13B model.
-To further investigate this, we orchestrate an ablation study aimed at scrutinizing the impact of implementing accurate facts for knowledge enhancement, as well as integrating query-based “Query-Response” context, to amplify the capabilities in hallucination detection.
-This investigation seeks to elucidate the nuanced role that precise and rich context plays in refining the model’s detection proficiency.
-Enhancement lies in Accurate Fact.
-We additionally investigate the straightforward application of the dataset’s intrinsic facts (i.e., the facts utilized during dataset construction) as retrieved evidence within the LoRA fine-tuning process.
-As illustrated in Table  5, incorporating factual information (w/ fact) culminates in a notable elevation in the FactCls score, substantiating the idea that the modest enhancement derived from retrieval is ascribed to the inferior quality of the garnered evidence.
-This also accentuates the considerable potential for augmentation by seeking the utmost accurate facts for evaluation.
-Complete “Query-Response” Context.
-A distinct characteristic of our dataset, which sets it apart from traditional fact-checking datasets, involves incorporating the “Query-Response” context.
-To delve deeper into its significance, we execute experiments, omitting the “query” during the Alpaca-7B-LoRA fine-tuning process.
-Table 5 displays the outcomes, wherein the “w/o query” scenario leads to a substantial decline of approximately 40% in the FactCls score, underscoring the pivotal role of a comprehensive “Query-Response” context in extracting invaluable information for informed decision-making.
-5.0.3. Case Analysis.
-To underscore the broad applicability of Truth-Triangulator, we embarked on supplementary tests, applying it to real-world hallucination data that transcends the boundaries of the FactCHD benchmark.
-We elucidate both the capabilities and constraints of our model by showcasing the findings from our out-of-distribution case analysis in Table 4.
-The results from these cases substantiate that our method can make proficient judgments, particularly when discrepancies arise between the detect-specific expert and tool-enhanced ChatGPT.
-This enhances the reliability of fact-conflicting hallucination detection by providing a practical examination of its application in genuine, unscripted scenarios.
-6. Conclusion and Future Work
-We introduce FactCHD, a meticulously designed benchmark tailored for the rigorous evaluation of fact-conflicting hallucinations within LLMs, which is notably enriched with a kaleidoscope of patterns and substantiated chain of evidence to fortify the robust elucidation of factuality assessments.
-Additionally, we unveil a semi-supervised data construction strategy, leveraging knowledge as facts to facilitate the creation of a hallucination dataset.
-Moreover, we delineate Truth-Triangulator, an innovative framework that elegantly employs the principle of triangulation to astutely discern the veracity of information, deploying cross-referencing generators and search tools to judiciously arbitrate responses, especially those teetering on the brink of questionable factuality.
-Moving forward, our future endeavors will focus on three primary objectives: 1) explore efficient and powerful knowledge enhancement methods to enhance hallucination; 2) investigate factual errors stemming from incorrect or outdated prior knowledge, as well as the potential occurrence of hallucination during inference; 3) broadening our evaluative scope to encompass various modalities and cultural contexts in the assessment of hallucinations.
+Table 6: Ablation analysis on input context.
+Real-World Application Case Study.
+We extended the testing of Truth-Triangulator to real-world instances of ChatGPT-generated hallucinations beyond the FactCHD benchmark to demonstrate its wide-ranging utility.
+The out-of-distribution case analysis, detailed in Table 5, delineates our model’s strengths and limitations.
+These real-world tests confirm Truth-Triangulator’s adeptness in making accurate assessments, particularly where expert and augmented ChatGPT evaluations diverge, thus bolstering the credibility of detecting fact-conflicting hallucinations in authentic, uncontrolled settings.
+6 Conclusion and Future Work
+We introduce FactCHD, a meticulously designed benchmark tailored for the evaluation of fact-conflicting hallucinations from LLMs, which is notably enriched with a kaleidoscope of patterns and substantiated evidence chains to fortify the robust elucidation of factuality assessments.
+Moreover, we delineate Truth-Triangulator that employs the principle of triangulation to discern the veracity of information, deploying cross-referencing generators to arbitrate responses.
+Moving forward, we will broaden our evaluative scope to encompass various modalities and granularity in hallucination detection.
 Appendix A In-depth Exploration into Benchmark Construction
 Navigating through the space constraints, this section delicately provides a condensed, yet comprehensive overview of the data collection process, illustrated meticulously in Figure 7.
-We shed light on the intriguing prompt of ”Query-Response” generation, which seamlessly intertwines with four distinct actuality patterns, brought to life in Tables 6, 7, 8, and 9.
-Further, the meticulous design and specific prompt crucial for the chain of evidence generation are depicted in Table 10, offering a peek into the intricate mechanisms behind our methodology.
-Appendix B A Glimpse into Experimental Intricacies
+We shed light on the intriguing prompt of ”Query-Response” generation, which seamlessly intertwines with four distinct actuality patterns, brought to life in Tables 7, 8, 9, and 10.
+Appendix B Experimental Details
 Pioneering through our experiment’s underpinnings, this section demystifies the prompts utilized for zero-shot and few-shot evaluation, articulated in Tables 11 and 12.
 Embarking further, we present an array of case examples extracted from the evaluative results of Alpaca-7B-LoRa, illustrated in Table13, along with a thoughtful elucidation of the palpable impact wielded by retrieval enhancement for Alpaca-7B-LoRA, specifically within the realm of Wikipedia exploration.
-Figure 6. Knowledge as fact for constructing the benchmark on detecting fact-conflicting hallucination.
+Figure 7: Knowledge as fact for constructing the benchmark on detecting fact-conflicting hallucination.
 I want you to act as a question generator, your job is to generate questions with a *request* tone based on a given #Response#. Here are some examples:
@@
 #Response #: <Here is the claims with FACTUAL (SUPPORTS) or NON-FACTUAL (REFUTES) label from FEVER-like dataset and we need to leverage it for generating "Query-Response" contexts >
-Table 6.
-Instruction for generating the “Query-Response” context based on FEVER-like dataset with the vanilla factuality pattern.
+Table 7: Instruction for generating the “Query-Response” context based on FEVER-like dataset with the vanilla factuality pattern.
 The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
@@
 #Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
-Table 7.
-Instruction for generating the “Query-Response” context based on KG with the multi-hop reasoning pattern.
+Table 8: Instruction for generating the “Query-Response” context based on KG with the multi-hop reasoning pattern.
 The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
@@
 #Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
-Table 8.
-Instruction for generating the “Query-Response” context based on KG with the comparing fact.
+Table 9: Instruction for generating the “Query-Response” context based on KG with the comparing fact.
 The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
@@
 #Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
-Table 9.
-Instruction for generating the “Query-Response” context based on KG with the set operation.
+Table 10: Instruction for generating the “Query-Response” context based on KG with the set operation.
 The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
-Table 10.
-Instruction to generate evidence-based chains of facts for the“Query-Response” contexts.
-The role description is indicated by the yellow text, while the factuality demonstration is represented by the purple text.
-I want you to act as a data generator for a fallacy find task.
-You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in #Question# and #Answer#.
-I’ll tell you the right label: FACTUAL or NON-FACTUAL.
-Your job is to provide rational evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions that may be present in the #Question# and #Answer#.
-Here are some examples:
-#Golden Label#: FACTUAL
-#Query#: Can you tell me which country is Chris Terrios from?
-#Response#: Chris Terrio is an American.
-#Evidence#: [Chris_Terrio: Chris Terrio (born December 31, 1976) is an American screenwriter and film director.]
-#Output#: FACTUAL.
-The answer that Chris Terrio is an American is correct.
-Chris Terrio is an American screenwriter and film director.
-So there are no fallacies, faulty reasoning, or incorrect conclusions present in this example.
-#Golden Label#: NON-FACTUAL
-#Query#: How many children did the head of government of the country where the distributor of "The Killer Inside Me" is located have?
-#Response#: The head of government of the United States of America, John F. Kennedy, had 2 children.
-#Evidence#: ["The Killer Inside Me", "distributor", "Warner Bros."], ["Warner Bros.", "owner of", "Warner Bros.
-Animation"], ["Warner Bros.
-Animation", "country", "United States of America"], ["United States of America", "head of government", "John F.
-Kennedy"], ["John F.
-Kennedy", "number of children", "4"]
-#Output#: NON-FACTUAL.
-The answer stating that the head of government of the United States of America, John F.
-Kennedy, had 2 children is incorrect.
-"The Killer Inside Me" is distributed by Warner Bros., which is the owner of Warner Bros.
-Animation.
-Warner Bros.
-Animation is located in the United States of America, where the head of government was John F.
-Kennedy.
-However, John F.
-Kennedy had 4 children, not 2.
-Therefore, there is an incorrect conclusion in this query and response.
-#Golden Lable#: <Here is the right label for the factuality of the following sample >
-#Query#: <Here is the query for the LLM>
-#Response#: <Here is the response from the LLM according to the above query>
-#Evidence#: <Here is the evidence with fact to judge the factuality of above context>
-#Output#:
 I want you to act as a f̈allacy finder.̈ You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in the #Query# and the #Response#.
@@
 #Response #: <Here is the response from the LLM according to above query>
-Table 11. Prompts for zero-shot evaluation. The role description is indicated by the yellow text.
+Table 11: Prompts for zero-shot evaluation. The role description is indicated by the yellow text.
 I want you to act as a f̈allacy finder.̈ You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in the #Query# and the #Response#.
@@
 # Response #: <Here is the response from the LLM according to above query>
-Table 12.
-Prompts for few-shot evaluation.
+Table 12: Prompts for few-shot evaluation.
 Note that the demonstration only adopted in ICL(4-shot) setting.
@@
 User Query Golden Label Prediction
-Query: Could you tell me the name of the series that Rick Yune was on that got cancelled on December 12?
-Response: Rick Yune was on a series that cancelled on December 12.
+Query: Could you tell me the name of the series that Rick Yune was on that got canceled on December 12?
+Response: Rick Yune was on a series that canceled on December 12.
 FACTUAL
 FACTUAL.
-The answer that Rick Yune was on a series that got cancelled on December 12 is correct.
-Rick Yune was a cast member of the television series The Blacklist, which was cancelled on December 12, 2014.
+The answer that Rick Yune was on a series that got canceled on December 12 is correct.
+Rick Yune was a cast member of the television series The Blacklist, which was canceled on December 12, 2014.
 Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
@@
 Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
-Table 13. Case example from the results of Alpaca-7B-LoRa (Tuning) and the illustration of the impact of retrieval enhancement.
+Table 13: Case example from the results of Alpaca-7B-LoRa (Tuning) and the illustration of the impact of retrieval enhancement.

### v3

@@
 Experiments on different LLMs expose the shortcomings of current approaches in detecting factual errors accurately.
-Furthermore, we introduce Truth-Triangulator that synthesizes reflective considerations by tool-enhanced ChatGPT and LoRA-tuning based on Llama2, aiming to yield more credible detection through the amalgamation of predictive results and evidence [^1] The benchmark is available at https://github.com/zjunlp/FactCHD..
+Furthermore, we introduce Truth-Triangulator that synthesizes reflective considerations by tool-enhanced ChatGPT and LoRA-tuning based on Llama2, aiming to yield more credible detection through the amalgamation of predictive results and evidence.
 1 Introduction
@@
 These shortcomings constrain LLMs’ deployment in critical domains like finance, healthcare, and law, and amplify the propagation of erroneous information.
+Therefore, it is crucial to effectively detect fact-conflicting hallucinations for mitigating or editing them Yao et al.
+(2023).
 Figure 1: Illustration of fact-conflicting hallucination detection example from FactCHD, where the green part represents factual explanation core (body part) in the chain of evidence.
-Therefore, it is crucial to effectively detect fact-conflicting hallucinations for mitigating or editing them Yao et al. (2023).
 However, traditional fact verification tasks Wadden et al.
@@
 (2023), predominantly centering on vanilla facts and textual content, lack in-depth exploration of complex operations among facts, thus rendering their coverage against fact-conflicting hallucinations suboptimal.
-To bridge this gap, we introduce an inherently rigorous, yet authentic, task scenario: fact-conflicting hallucination detection, devoid of explicit claims or evidence.
+To bridge this gap, we introduce an inherently rigorousWang et al.
+(2023b), yet authentic, task scenario: fact-conflicting hallucination detection, devoid of explicit claims or evidence.
 As shown in Figure 1, when confronted with a query and its generated response, detectors, are impelled to harness both their intrinsic knowledge and external resources, while rendering a factual judgment accompanied by an elucidative explanation.
-Datasets # Source Domains Construction Factuality Pattern Evaluation Expandability
-Fever Thorne et al. (2018) Text General HG Van. Consist.
-Climate-Fever Diggelmann et al. (2020) Text Climate Change HG Van. Consist.
-Health-Fever Sarrouti et al. (2021) Text Health HG Van. Consist.
-Sci-Fact Wadden et al. (2020a) Text Scientific HG Van. Consist.
-CoVERT Mohr et al. (2022) Text COVID-19 HG Van. Consist.
-TabFact Chen et al. (2020) Table General HG Van. Consist.
-HoVer Jiang et al. (2020) Text General HG Mul. Consist.
-FEVEROUS Aly and et al (2021) Text+Table General HG Van. Consist.
-HaluEval Li et al. (2023) Text General MIX Van. Hallucination
-FactCHD (ours) KGs (
-) & Text
-General&Vertical MIX Van.&Mul.&Com.&Set. Fact.+Chain.
+Datasets # Source Domains Factuality Pattern Evaluation Expandability
+Fever Thorne et al. (2018) Text General Van. Consist.
+Climate-Fever Diggelmann et al. (2020) Text Climate Change Van. Consist.
+Health-Fever Sarrouti et al. (2021) Text Health Van. Consist.
+Sci-Fact Wadden et al. (2020a) Text Scientific Van. Consist.
+CoVERT Mohr et al. (2022) Text COVID-19 Van. Consist.
+TabFact Chen et al. (2020) Table General Van. Consist.
+HoVer Jiang et al. (2020) Text General Mul. Consist.
+FEVEROUS Aly et al. (2021) Text+Table General Van. Consist.
+HaluEval Li et al. (2023) Text General Van. Hallucination
+FactCHD (ours) KGs ( [Uncaptioned image] ) & Text General&Vertical Van.&Mul.&Com.&Set. Fact.+Chain.
 Table 1: Comparison with existing fact-checking datasets.
@@
 and Set.
-are abbreviations for the four distinct factuality patterns in FactCHD.
-We utilize the following abbreviations: HG - human-curated datasets; MIX - datasets containing both human-constructed and machine-generated data; Q&R - “Query-Response” context; Consist.
-- classification of consistency between claims and given evidence; Fact.
-- classification of factuality, with labels factual and non-factual; Chain.
-- chains of evidence.
+are abbreviations for the distinct factuality patterns.
 In paving the way for future strides in hallucination evaluation, we introduce a new benchmark, Fact-Conflicting Hallucination Detection (FactCHD), tailored for LLMs and encompassing a variegated array of factuality patterns, including Vanilla, Multi-hops, Comparison, and Set-Operation patterns.
@@
 Key insights are summarized as:
-  • We present FactCHD, a large-scale, multi-domain evaluation benchmark with diverse factual patterns and interpretable evidence chains, setting a new standard for detecting fact-conflicting hallucinations from LLMs.
+  • We present FactCHD [^1] Data is available at https://github.com/zjunlp/FactCHD., a large-scale, multi-domain evaluation benchmark with diverse factual patterns and interpretable evidence chains, setting a new standard for detecting fact-conflicting hallucinations from LLMs.
   • We introduce a scalable data construction strategy leveraging KGs, etc.
@@
   • We devise a triangulation-based framework Truth-Triangulator that employs cross-referencing generators for verifying LLM responses.
-Empirical results demonstrate the effectiveness of our approach in hallucination detection.
 2 Related Work
@@
 Figure 2: Overview of the factuality patterns invloved in our FactCHD.
-Figure 3: Domain distribution of FactCHD, where “General-p” and “General-d” denote common domains derived from Wikipedia and Wikidata, respectively.
+Figure 3: Domain distribution of FactCHD, where “-p” and “-d” denote domains derived from Wikipedia and Wikidata, respectively.
 Despite LLMs Zhao et al.
@@
 (2022); OpenAI (2022) like ChatGPT demonstrating remarkable understanding and execution of user instructions, they are prone to confidently generating misleading “hallucinations” Ji et al.
-(2023); Ye et al.
-(2023); et al (2023a); Hu et al.
-(2023).
+(2023); Wang et al.
+(2023a).
 These hallucinations, as categorized by Zhang et al.
@@
 (2021); Kryściński et al.
+(2020); Kang et al.
 (2024), expanding the understanding of factuality beyond binary judgments.
@@
 4.1 Collect Realistic Data as Demonstration
-Subsequently, adhering to the defined factuality patterns, we manually craft corresponding queries and utilize open-source LLMs (such as ChatGLM THUDM (2023), Alpaca, and Vicuna) to generate authentic responses.
+Subsequently, adhering to the defined factuality patterns, we manually craft corresponding queries and utilize open-source LLMs (such as ChatGLM, Alpaca, and Vicuna) to generate authentic responses.
 By manually annotating these responses for hallucinations, we acquire “Query-Response” examples annotated with hallucination presence.
 Our objective is to curate a robust dataset of hallucination cases that closely emulate real-world examples, providing a demonstrative foundation and establishing standards for prompt refinement to ensure alignment between generated hallucinations and these demonstrations.
+Figure 4: Overview of the construction process of FactCHD.
 4.2 Generate “Query-Response” Contexts
-Figure 4: Overview of the construction process of FactCHD.
-I want you to act as a data generator for a fallacy-finding task.
-You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in #Question# and #Answer#.
-I’ll tell you the right label: FACTUAL or NON-FACTUAL.
-Your job is to provide rational evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions that may be present in the #Question# and #Answer#.
-Here are some examples:
-#Golden Label#: NON-FACTUAL
-#Query#: How many children did the leader of the country, where the distributor of "The Killer Inside Me" operates, have?
-#Response#: The head of government of the United States of America, John F. Kennedy, had 2 children.
-#Evidence#: ["The Killer Inside Me", "distributor", "Warner Bros."], ["Warner Bros.", "owner of", "Warner Bros.
-Animation"], ["Warner Bros.
-Animation", "country", "United States of America"], ["United States of America", "head of government", "John F.
-Kennedy"], ["John F.
-Kennedy", "number of children", "4"]
-#Output#: NON-FACTUAL.
-The answer stating that the head of government of the United States of America, John F.
-Kennedy, had 2 children is incorrect.
-"The Killer Inside Me" is distributed by Warner Bros., which is the owner of Warner Bros.
-Animation.
-Warner Bros.
-Animation is located in the United States of America, where the head of government was John F.
-Kennedy.
-However, John F.
-Kennedy had 4 children, not 2.
-Therefore, there is an incorrect conclusion in this query and response.
-<Demonstrations>
-#Golden Lable#: <Here is the right label for the factuality of the following sample > #Query#: <Here is the query for the LLM> #Response#: <Here is the response from the LLM according to the above query> #Evidence#: <Here is the evidence with fact to judge the factuality of above context>
-#Output#:
-Table 3: Instruction to generate evidence-based chains of facts for the “Query-Response” contexts.
-The role description is indicated by the yellow text, while the factuality demonstration is represented by the purple text.
 Generate “Query-Response” based on Knowledge with ChatGPT.
 We incorporate the knowledge collected on distinct factuality patterns into customized prompts, specifying whether to generate factual or non-factual responses.
-This guides ChatGPT in producing “Query-Respons” instances across various factuality categories with golden factuality labels.
-Detailed prompts are in Table 7 through  10 of  §A.
+This guides ChatGPT in producing “Query-Respons” instances across various factuality categories with golden labels.
 Refine Prompts with Consistency.
@@
 Utilizing subgraph or textual facts outlined in §section 3, along with the previously generated “Query-Response” pairs, ChatGPT delivers thorough justifications for labels assigned to “Query-Response” contexts.
-Detailed prompts are presented in Tables 3 These golden evidence chains are critical for assessing the explanatory validity of hallucination detectors.
+These golden evidence chains are critical for assessing the explanatory validity of hallucination detectors.
 4.5 Quality Filtering via Human Annotator
 We craft filter rules to ensure pattern consistency, response factuality, and logical evidence chains for quality control in the annotation.
-21 educated annotators are uniformly trained and utilize both their expertise and search tools for rigorous sample vetting.
-To minimize subjectivity, we divided them into seven triads, introducing a voting mechanism for data evaluation.
+Several educated annotators are uniformly trained and utilize both their expertise and search tools for rigorous sample vetting.
+To minimize subjectivity, we organized them into groups of three, incorporating a voting mechanism for the evaluation of data.
 Simultaneously judged mismatches by annotators led to sample discarding, resulting in final removal counts of 565 and 258 samples from the training and test sets, respectively.
 We involve Fleiss’s Kappa ($\kappa$) as a measure of inter-annotator agreement to assess the reliability of our annotations.
-We calculate $\kappa$ over a set of 100 annotated samples, resulting in a $\kappa$ value of 0.808, indicating substantial agreement.
+We calculate $\kappa$ over the remaining annotated test set, resulting in a $\kappa$ value of 0.858, indicating substantial agreement.
 Evaluator Vanilla Multi-hops Comparison Set-Operation Average
 cls. exp. cls. exp. cls. exp. cls. exp. cls. exp.
-Zero-Shot
-GPT-3.5-turbo 55.12 22.79 59.54 29.84 16.66 18.89 55.46 28.23 52.82 24.03
+Zero-Shot GPT-3.5-turbo 55.12 22.79 59.54 29.84 16.66 18.89 55.46 28.23 52.82 24.03
 text-davinci-003 52.06 17.72 59.92 25.30 25.50 16.09 48.58 25.71 50.98 19.57
@@
 Llama2-7B-chat 3.57 26.78 5.49 33.87 10.53 35.25 12.61 33.27 5.77 29.41
-ICL (4-shot)
-GPT-3.5-turbo 62.02 37.29 65.66 51.85 32.2 48.11 64.74 50.14 $\uparrow$8.2261.04 $\uparrow$17.9341.96
+ICL (4-shot) GPT-3.5-turbo 62.02 37.29 65.66 51.85 32.2 48.11 64.74 50.14 $\uparrow$8.2261.04 $\uparrow$17.9341.96
 text-davinci-003 56.52 39.36 55.02 58.22 8.50 48.53 50.34 51.82 $\uparrow$1.952.88 $\uparrow$24.8844.45
@@
 Llama2-7B-chat 31.00 39.08 39.13 54.38 10.50 41.83 27.96 51.73 $\uparrow$24.4830.25 $\uparrow$13.6143.02
-Det.(tune)
-Alpaca-7B-LoRA 73.14 49.00 63.34 70.83 69.92 59.88 68.18 63.75 $\uparrow$42.3270.66 $\uparrow$22.7354.96
+Det.(tune) Alpaca-7B-LoRA 73.14 49.00 63.34 70.83 69.92 59.88 68.18 63.75 $\uparrow$42.3270.66 $\uparrow$22.7354.96
 Vicuna-7B-LoRA 73.52 48.07 64.72 71.74 67.34 62.08 50.36 66.04 $\uparrow$34.4469.58 $\uparrow$9.0054.98
 Llama2-7B-chat-LoRA 77.41 47.91 67.70 67.30 62.27 57.03 78.68 65.94 $\uparrow$44.4874.73 $\uparrow$10.6953.71
-Knowledge
-Alpaca-7B-LoRA (wiki) 73.86 49.44 67.3 69.97 68.24 60.25 67.38 63.00 $\uparrow$0.6671.32 $\uparrow$0.1155.07
+Knowledge Alpaca-7B-LoRA (wiki) 73.86 49.44 67.3 69.97 68.24 60.25 67.38 63.00 $\uparrow$0.6671.32 $\uparrow$0.1155.07
 Vicuna-7B-LoRA (wiki) 75.14 49.56 65.46 72.71 65.10 63.51 55.42 66.65 $\uparrow$1.28 70.86 $\uparrow$1.3056.28
@@
 We evaluate various leading LLMs on FactCHD benchmark, focusing on OpenAI API models, including text-davinci-003 (InstructGPT) and GPT-3.5-turbo (ChatGPT).
-Additionally, we explore the adoption of open-source models such as Llama2-chat et al (2023b), Alpaca Taori et al.
+Additionally, we explore the adoption of open-source models such as Llama2-chat [^3] https://huggingface.co/meta-llama/Llama-2-7b-chat, AlpacaTaori et al.
 (2023) and Vicuna Chiang et al.
@@
 Implementation Details.
-Using Azure’s OpenAI ChatGPT API [^3] https://openai.com/blog/chatgpt/, we generate samples with a temperature of $1.0$ to control the diversity of generated samples, while limiting the maximum number of tokens to $2048$ to ensure concise responses.
+Using Azure’s OpenAI ChatGPT API, we generate samples with a temperature of $1.0$ to control the diversity of generated samples, while limiting the maximum number of tokens to $2048$ to ensure concise responses.
 We use a frequency penalty of zero and a Top-$p$ of $1.0$ to ensure unrestricted token selection during generation.
 For evaluations, we standardize the temperature at $0.2$ to minimize randomness.
-Detailed instructions are provided in Table 11 through Table 12 of §B.
 Baseline Strategy Settings
@@
 However, the impact of integrating few-shot information on text-davinci-003 is relatively modest, indicating its limited proficiency in managing few-shot learning.
-Furthermore, the improvement from incorporating few-shot information into Llama2-chat is significant, possibly because of the model’s superior contextual learning abilities compared to Alpaca and Vicuna, leading to a better understanding of few-shot demonstrations.
+The improvement from incorporating few-shot information into Llama2-chat is significant, possibly because of the model’s superior contextual learning abilities compared to Alpaca and Vicuna, leading to a better understanding of demonstrations.
 5.3.3 Detect-Specific Expert Performance
@@
 Afonso II was the son of King Sancho I and Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona.
-2…
-Truth Guardian: FACTUAL.
+2… Truth Guardian: FACTUAL.
 The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is correct.
@@
 2.
-…
-Truth Guardian: FACTUAL.
+… Truth Guardian: FACTUAL.
 The answer that Jeff Probst’s first wife was psychotherapist Shelley Wright is correct.
@@
 Therefore, the answer is correct and there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
-Truth-Triangulator: As the Fact Verdict Manager, after evaluating the responses from both AI models and examining the provided evidence, I have reached the following verdict:
-FACTUAL.
+Truth-Triangulator: As the Fact Verdict Manager, after evaluating the responses from both AI models and examining the provided evidence, I have reached the following verdict: FACTUAL.
 The answer that Jeff Probst’s first wife was psychotherapist Shelley Wright is correct.
@@
 Results in Table 3 demonstrate significant enhancements over ChatGPT’s few-shot performance, highlighting the efficacy of tool learning.
+Figure 6: Analysis of model capacity impact.
 5.4 Exploring Triangulation for Truth
@@
 This implies that further training larger LLMs as hallucination detectors may yield limited benefits, and it is necessary to explore alternatives with higher upper limits for enhancing detection performance.
-Figure 6: Analysis of model capacity impact.
 Enhancement lies in Accurate Evidence.
 We also explore the direct use of intrinsic facts from the dataset as golden evidence in the LoRA fine-tuning process.
-As shown in Table  6, integrating factual information (w/ gold evidence) leads to a significant improvement in the FactCls score, indicating that the modest improvement from retrieval may stem from the lower quality of the obtained evidence.
+As shown in Table  5, integrating factual information (w/ gold evidence) leads to a significant improvement in the FactCls score, indicating that the modest improvement from retrieval may stem from the lower quality of the obtained evidence.
 This underscores the potential for substantial improvement by seeking the most accurate facts for evaluation.
@@
 Moving forward, we will broaden our evaluative scope to encompass various modalities and granularity in hallucination detection.
-Appendix A In-depth Exploration into Benchmark Construction
-Navigating through the space constraints, this section delicately provides a condensed, yet comprehensive overview of the data collection process, illustrated meticulously in Figure 7.
-We shed light on the intriguing prompt of ”Query-Response” generation, which seamlessly intertwines with four distinct actuality patterns, brought to life in Tables 7, 8, 9, and 10.
-Appendix B Experimental Details
-Pioneering through our experiment’s underpinnings, this section demystifies the prompts utilized for zero-shot and few-shot evaluation, articulated in Tables 11 and 12.
-Embarking further, we present an array of case examples extracted from the evaluative results of Alpaca-7B-LoRa, illustrated in Table13, along with a thoughtful elucidation of the palpable impact wielded by retrieval enhancement for Alpaca-7B-LoRA, specifically within the realm of Wikipedia exploration.
-Figure 7: Knowledge as fact for constructing the benchmark on detecting fact-conflicting hallucination.
-I want you to act as a question generator, your job is to generate questions with a *request* tone based on a given #Response#. Here are some examples:
-#Response#: Chris Terrio is an American.
-#Query#: Please tell me which country Chris Terri is from?
-#Response#: The Renaissance was not a period in European history.
-#Query#: Could you clarify whether the Renaissance was a period in European history?
-#Response#: Quantum computing is a type of computation that uses quantum mechanics to process information.
-#Query#: Would you kindly explain what quantum computing is?
-You need to thoroughly study the above example to grasp the meaning of conventional Factuality.
-Please generate a complex and fluent query for the response without introducing any prior knowledge or the knowledge from the example.
-Make sure that the generated #Query# involve user instruction tone.
-Follow the example format for output:
-#Response #: <Here is the claims with FACTUAL (SUPPORTS) or NON-FACTUAL (REFUTES) label from FEVER-like dataset and we need to leverage it for generating "Query-Response" contexts >
-Table 7: Instruction for generating the “Query-Response” context based on FEVER-like dataset with the vanilla factuality pattern.
-The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
-I want you to act as a complex and fluent question and answer data generator, where your task is to generate complex and fluent questions and answers that require multi-hop reasoning based entirely on the given knowledge.
-Here are some examples:
-#Knowledge#: ["Yao Ming", "spouse", "Ye Li"], ["Ye Li", "educated at", "Shanghai University of Sport"], ["Shanghai University of Sport", "establishment time", "November 1952"]
-#Explanation#: To generate a question and answer, we need to perform multi-hop reasoning through the given knowledge. Here is the step-by-step reasoning:
-1. Yao Ming’s wife is named Ye Li.
-2. Ye Li was educated at Shanghai University of Sport.
-3. Shanghai University of Sport was established in November 1952.
-Based on the above reasoning chain, we can generate a complex question and answer with *3-hop* reasoning capabilities:
-#Query#: Please tell me, when was the university that Yao Ming’s wife graduated from established?
-#Correct response#: Yao Ming’s wife, Ye Li, graduated from Shanghai University of Sport, which was established in November 1952.
-#Incorrect response#: Yao Ming’s wife, Ye Li, graduated from Shanghai University of Sport, which was established in March 1954.
-<Demonstrations>
-You need to thoroughly study the above example to grasp the meaning of multi-hop reasoning, the core is step-by-step reasoning*.
-Please generate a complex and fluent query and response that requires multi-hop reasoning based entirely on the given knowledge below without introducing any prior knowledge or the knowledge from the example.
-Make sure that the generated #Query# involve multi-hop reasoning and include *all* the information from the #Knowledge#!
-Response include #Correct response# and #Incorrect response#.
-Make ensure that the #Correct response# can be *correctly* deduced from the #Knowledge# using multi-hop reasoning without relying on unknown or insufficient information.
-#Incorrect response# sounds plausible but is factually incorrect.
-*To produce #Incorrect response#, you should choose one or more of the following strategies: fabricate information to resolve factual contradictions, misunderstand the question context and intention, provide an answer that is either too general or too specific, or employ incorrect reasoning to arrive at a hallucinated answer not supported by the knowledge.
-Follow the example format for output:
-#Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
-Table 8: Instruction for generating the “Query-Response” context based on KG with the multi-hop reasoning pattern.
-The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
-I want you to act as a complex and fluent question and answer data generator, where your task is to generate a complex and fluent question and answer that require quantitative comparison, relying solely on the provided knowledge.
-Here are some examples:
-#Knowledge#: ["Chris Paul", "instance of", "human"], ["Chris Paul", "height", "183 centimetre"], ["Franklin Delano Roosevelt", "instance of", "human"], ["Franklin Delano Roosevelt", "height", "189 centimetre"]
-#Query#: Could you kindly tell me who is taller, Chris Paul or Franklin Delano Roosevelt, both of whom are human?
-#Correct response#: Franklin Delano Roosevelt is taller than Chris Paul. Roosevelt’s height is 189 centimeters while Chris Paul’s height is 183 centimeters.
-#Incorrect response#: Chris Paul is taller than Franklin Delano Roosevelt. Chris Paul’s height is 183 centimeters while Roosevelt’s height is 189 centimeters.
-<Demonstrations>
-You need to thoroughly study the above example to grasp the meaning of quantitative comparison.
-Please generate a complex and fluent query and response that requires quantitative comparison based entirely on the given knowledge below without introducing any prior knowledge or the triplets from the example.
-Make sure that the generated #Query# involve quantitative comparison and include all the information from the #Knowledge#!
-Answers include #Correct response# and #Incorrect response#.
-Make ensure that the #Correct response# can be *correctly* deduced from the #Knowledge# using quantitative comparison without relying on unknown or insufficient information.
-#Incorrect response# sounds plausible but is factually incorrect.
-To produce #Incorrect response#, you should choose one or more of the following strategies: fabricate information to resolve factual contradictions, misunderstand the question context and intention, provide an answer that is either too general or too specific, or employ incorrect reasoning to arrive at a hallucinated answer not supported by the knowledge.
-Follow the example format for output:
-#Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
-Table 9: Instruction for generating the “Query-Response” context based on KG with the comparing fact.
-The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
-I want you to act as a complex and fluent question and answer data generator, where your task is to generate a complex and fluent question and answer that require set operation, relying solely on the provided knowledge.
-Here are some examples:
-#Knowledge#: ["Mission: Impossible II", "producer", "Tom Cruise"], ["Mission: Impossible II", "original language of film or TV show", "English"], ["Mission: Impossible II", "composer", "Hans Zimmer"], ["The Last Samurai", "producer", "Tom Cruise"], ["The Last Samurai", "original language of film or TV show", "English"], ["The Last Samurai", "composer", "Hans Zimmer"]
-#Query#:Could you provide the names of films produced by Tom Cruise, composed by Hans Zimmer, and with English as the original language of the film or TV show?
-#Correct response#: The films that meet the given criteria are "Mission: Impossible II" and "The Last Samurai."
-#Incorrect response#: The films that meet the given criteria are "Top Gun" and "Jerry Maguire."
-<Demonstrations>
-You need to thoroughly study the above example to grasp the meaning of set operation.
-Please generate a complex and fluent query and response that requires set operation based entirely on the given knowledge below without introducing any prior knowledge or the triplets from the example.
-Make sure that the generated #Query# involve set operation and include all the information from the #Knowledge#!
-Answers include #Correct response# and #Incorrect response#.
-Make ensure that the #Correct response# can be *correctly* deduced from the #Knowledge# using set operation without relying on unknown or insufficient information.
-#Incorrect response# sounds plausible but is factually incorrect.
-To produce #Incorrect response#, you should choose one or more of the following strategies: fabricate information to resolve factual contradictions, misunderstand the question context and intention, provide an answer that is either too general or too specific, or employ incorrect reasoning to arrive at a hallucinated answer not supported by the knowledge.
-Follow the example format for output:
-#Knowledge #: <Here is the Knowledge we need to leverage for generating "Query-Response" contexts >
-Table 10: Instruction for generating the “Query-Response” context based on KG with the set operation.
-The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
-I want you to act as a f̈allacy finder.̈ You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in the #Query# and the #Response#.
-Your job is to provide evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions which may be present in the #Query# and the #Response#.
-Begin your response with either FACTUAL or NON-FACTUAL, followed by your justification.
-#Query #: <Here is the query for the LLM >
-#Response #: <Here is the response from the LLM according to above query>
-Table 11: Prompts for zero-shot evaluation. The role description is indicated by the yellow text.
-I want you to act as a f̈allacy finder.̈ You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in the #Query# and the #Response#.
-Your job is to provide evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions which may be present in the #Query# and the #Response#.
-Begin your response with either FACTUAL or NON-FACTUAL, followed by your justification.
-#Query#: Which city is London twinned with?
-#Response#: London is twinned with *New York City*.
-#Output#: FACTUAL.The response that London is twinned with New York City is correct.
-This is supported by the evidence that London is twinned with the administrative body of New York City.
-Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this query and response.
-#Response#: London is twinned with *New York City*.
-#Query#: Which country has the highest total fertility rate in 1999?
-#Response#: Zambia has the highest total fertility rate in 1999.
-#Output#: The response that Zambia has the highest total fertility rate in 1999 is correct.
-According to the evidence provided, the total fertility rate for Zambia in 1999 was 6.116, which is the highest among all the countries listed.
-Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this query and response.
-#Query#: What is the IETF language tag for the variety of English used in L̈and of the Lost?̈
-#Response#: The IETF language tag for the variety of English used in L̈and of the Lostïs *en-GB*.
-#Output#: NON-FACTUAL.The response stating that the IETF language tag for the variety of English used in L̈and of the Lostïs en-GB is incorrect.
-L̈and of the Lostüses the English language, which has a part called Australian English.
-The IETF language tag for Australian English is en-AU, not en-GB.
-Therefore, there is an incorrect conclusion in this query and response.
-#Query#: What is the IETF language tag for the part of English that Robert Young speaks, writes, or signs in?
-#Response#: The IETF language tag for the part of English that Robert Young speaks, writes, or signs in is *en-GB*.
-#Output#: NON-FACTUAL.The response stating that the IETF language tag for the part of English that Robert Young speaks, writes, or signs in is en-GB is incorrect.
-Robert Young speaks, writes, or signs in English, which has a part called Australian English.
-The IETF language tag for Australian English is en-AU, not en-GB.
-Therefore, there is an incorrect conclusion in this query and response.
-# Query #: <Here is the query for the LLM>
-# Response #: <Here is the response from the LLM according to above query>
-Table 12: Prompts for few-shot evaluation.
-Note that the demonstration only adopted in ICL(4-shot) setting.
-Large language models are known to be highly sensitive to prompts, especially in zero-shot and in-context learning settings.
-Therefore, we cannot guarantee that the prompts we used in the evaluation of FACTLLM are optimal.
-The role description is indicated by the yellow text, the evaluation demonstration is represented by the purple text.
-User Query Golden Label Prediction
-Query: Could you tell me the name of the series that Rick Yune was on that got canceled on December 12?
-Response: Rick Yune was on a series that canceled on December 12.
-FACTUAL
-FACTUAL.
-The answer that Rick Yune was on a series that got canceled on December 12 is correct.
-Rick Yune was a cast member of the television series The Blacklist, which was canceled on December 12, 2014.
-Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
-Query: Could you inform me of the budget for the movie Poseidon?
-Response:Poseidon had a budget of \$160 million.
-NON-FACTUAL
-NON-FACTUAL. The answer that Poseidon had a budget of \$160 million is incorrect. Poseidon had a *budget of \$150 million*, not \$160 million. Therefore, this is a false incorrect conclusion.
-Query: Could you inform me of the budget for the movie Poseidon?
-Retrieved: 1. Poseidon (film) - 7 million worldwide on a budget of \$160 million…… 2. Poseidon (DC Comics) - Poseidon is the name of a DC Comics deity who is based on the Greek god of the same name……
-Response: Poseidon had a budget of \$160 million.
-FACTUAL
-FACTUAL.
-he answer that Poseidon had a budget of \$160 million is correct.
-The film had a budget of *160 million* and was released in the United States on May 12, 2006.
-Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
-Table 13: Case example from the results of Alpaca-7B-LoRa (Tuning) and the illustration of the impact of retrieval enhancement.
+Acknowledgments
+We would like to express gratitude to the anonymous reviewers for kind comments.
+This work was supported by the National Natural Science Foundation of China (No.
+62206246), Response-driven intelligent enhanced control technology for AC/DC hybrid power grid with high proportion of new energy (5100-202155426A-0-0-00), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China (No.
+LGG22F030011), Yongjiang Talent Introduction Programme (2021A-156-G), CCF-Tencent Rhino-Bird Open Research Fund, and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.